### v1

# Irreducibility of three dimensional automorphic Galois representations over totally real fields

Abstract.
Let $\rho_{\ell}$ be a semisimple $\ell$-adic representation of a number field $K$ that is unramified almost everywhere. We introduce a new notion called weak abelian direct summands of $\rho_{\ell}$ and completely characterize them, for example, if the algebraic monodromy of $\rho_{\ell}$ is connected. If $\rho_{\ell}$ is in addition $E$-rational for some number field $E$, we prove that the weak abelian direct summands are locally algebraic (and thus de Rham). We also show that the weak abelian parts of a connected semisimple Serre compatible system form again such a system.
Using our results on weak abelian direct summands, when $K$ is totally real and $\rho_{\ell}$ is the three-dimensional $\ell$-adic representation attached to a regular algebraic cuspidal automorphic, not necessarily polarizable representation $\pi$ of $\mathrm{GL}_{3}(\mathbb{A}_{K})$ together with an isomorphism $\mathbb{C}\simeq\overline{\mathbb{Q}}_{\ell}$, we prove that $\rho_{\ell}$ is irreducible. We deduce in this case also some $\ell$-adic Hodge theoretic properties of $\rho_{\ell}$ if $\ell$ belongs to a Dirichlet density one set of primes.
Mathematics Subject Classification (2010): 11F80, 11F70, 11F22, 20G05.
1. Introduction
Let $K$ be a totally real (or CM) number field, $\pi$ a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{n}(\mathbb{A}_{K})$, and $\iota:\mathbb{C}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\overline{\mathbb{Q}}_{\ell}$ a field isomorphism. Denote by $\operatorname{Gal}_{K}$ the absolute Galois group of $K$, by $\Sigma_{K}$ the set of finite places of $K$, by $S_{\ell}\subset\Sigma_{K}$ the subset of places $v$ that divide $\ell$, and by $S\subset\Sigma_{K}$ the subset of places $v$ such that the local component $\pi_{v}$ is ramified. Attached to the pair $(\pi,\iota)$ is a semisimple (conjecturally irreducible, see e.g., [Ram08]) $\ell$-adic representation of $K$
(1) $\rho_{\pi,\iota}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$
that satisfies local-global compatibility away from $\ell$: if $v$ is a finite place of $K$ outside $S_{\ell}$, then
(2) $\mathrm{WD}(\rho_{\pi,\iota}|_{\operatorname{Gal}_{K_{v}}})^{ss}=\iota\circ\mathrm{rec}_{K_{v}}(\pi_{v}\otimes|\det|_{v}^{\frac{1-n}{2}}).$
The construction of $\rho_{\pi,\iota}$ such that (2) holds for all $v$ outside $S\cup S_{\ell}$ was given by [HLTT16] (and resp. [Sc15]) and the validity of (2) for all $v$ outside $S_{\ell}$ was obtained in [Va18]. If $\pi$ is in addition polarized, the construction of such an $\ell$-adic representation was done earlier and moreover, the following $\ell$-adic Hodge properties hold for all $v\in S_{\ell}$:
  1. (dR):
    $\rho_{\pi,\iota}|_{\operatorname{Gal}_{K_{v}}}$ is de Rham with determined Hodge-Tate numbers;
  2. (Cr):
    $\rho_{\pi,\iota}|_{\operatorname{Gal}_{K_{v}}}$ is crystalline if $\pi_{v}$ is unramified,
see [BLGGT14b, §$\mathsection 2.1$] and the references therein for more details. These properties are crucial to investigating the irreducibility (resp. residual irreducibility if $\ell\gg 0$) of (1) [CG13],[PT15],[Xi19],[Hu23a] by using potential automorphy theorems [BLGGT14b] (resp. the method of algebraic envelopes [Hu23a]). For general $\pi$, it is not even known that (1) is Hodge-Tate at $v\in S_{\ell}$ unless more conditions (e.g., residual irreducibility of $\rho_{\pi,\iota}$) are assumed, see [ACC+23],[A’C23].
When $K$ is totally real and $n=3$ (resp. $4$), D. Ramakrishnan obtained (by using an L-function argument) the irreducibility of (1) assuming some $\ell$-adic Hodge properties [Ram13, Theorem B]. In this article, we prove such irreducibility unconditionally (Theorem 1.1) and deduce in return, the properties (dR) and (Cr) above, if $\rho_{\ell}$ is of irreducible type ($A_{2}$) and $\ell$ belongs to a Dirichlet density one set of rational primes (Theorem 3.6).
Theorem 1.1.
Let $K$ be a totally real field, $\pi$ a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{3}(\mathbb{A}_{K})$, and $\iota:\mathbb{C}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\overline{\mathbb{Q}}_{\ell}$ a field isomorphism. Then the three-dimensional $\ell$-adic Galois representation (1) of $K$ attached to $(\pi,\iota)$ is irreducible.
The new input for Theorem 1.1 is the following result which is of independent interest.
Theorem 1.2.
Let $K$ and $E\subset\overline{\mathbb{Q}}_{\ell}$ be number fields, and $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ a semisimple $E$-rational $\ell$-adic representation of $K$. If $\psi_{\ell}$ is a weak abelian direct summand of $\rho_{\ell}$, then it is locally algebraic, and in particular de Rham.
We briefly describe the remaining sections of this article. We first recall basic notions and facts about (abelian) $\ell$-adic representations in §$\mathsection 2.1$§$\mathsection 2.5$. Then we introduce weak abelian direct summands (a notion weaker than abelian subrepresentations) of an almost everywhere unramified semisimple $\ell$-adic Galois representation and completely characterize them in case the algebraic monodromy is connected (Proposition 2.9) in §$\mathsection 2.6$. Theorem 1.2 is established in §$\mathsection 2.7$ by using §$\mathsection 2.1$§$\mathsection 2.6$ and a remarkable result of Waldschmidt. Finally in §$\mathsection 2.8$§$\mathsection 2.9$, we formulate a conjecture on $E$-rationality of the abelian part $\rho_{\ell}^{\text{ab}}$ (Conjecture 2.14) with some evidence (Theorem 2.15) and prove the compatibility of the weak abelian part of a connected semisimple Serre compatible system (Theorem 2.20).
[…]
2. Weak abelian direct summands of $\ell$-adic representation
We fix some notation. Let $K$ be a number field and $\Sigma_{K}$ the set of finite places of $K$. For a rational prime $\ell$, denote by $S_{\ell}$ the set of finite places $v\in\Sigma_{K}$ above $\ell$. An $\ell$-adic representation of $K$ is a continuous homomorphism $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$. The algebraic monodromy group $\mathbf{G}_{\ell}$ of $\rho_{\ell}$ is the Zariski closure of the image $\rho_{\ell}(\operatorname{Gal}_{K})$ in $\mathrm{GL}_{n,\overline{\mathbb{Q}}_{\ell}}$.
2.1. $E$-rationality
Let $E\subset\overline{\mathbb{Q}}_{\ell}$ be a number field. An $n$-dimensional $\ell$-adic representation
$\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$
is said to be $E$-rational [Se98, Chap. I] if there is a finite subset $S\subset\Sigma_{K}$ such that for every finite place $v\in\Sigma_{K}\backslash S$, the representation $\rho_{\ell}$ is unramified at $v$ and the characteristic polynomial of $\rho_{\ell}(\operatorname{Frob}_{v})$ has coefficients in $E$, i.e.,
$\det(\rho_{\ell}(\operatorname{Frob}_{v})-T\cdot\text{Id})\in E[T],$
where $\operatorname{Frob}_{v}$ denotes the Frobenius class at $v$.
2.2. Serre compatible system
Let $S\subset\Sigma_{K}$ a finite subset and $E$ a number field. For $\lambda\in\Sigma_{E}$, denote its residue characteristic by $\ell(\lambda)$ (and so that $\overline{E}_{\lambda}\cong\overline{\mathbb{Q}}_{\ell(\lambda)}$). A family of $n$-dimensional $\ell$-adic representations (indexed by $\Sigma_{E}$)
(3) $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda}):~{}\lambda\in\Sigma_{E}\}$
is said to be a Serre compatible system (SCS) of $K$ defined over $E$ unramified outside $S$ ([Se98, Chap. 1]) if there exists $P_{v}(T)\in E[T]$ for all $v\in\Sigma_{K}\backslash S$ such that for any $\lambda\in\Sigma_{E}$ and $v\in\Sigma_{K}\backslash(S\cup S_{\ell(\lambda)})$, the representation $\rho_{\lambda}$ is unramified at $v$ and
$\det(\rho_{\lambda}(\operatorname{Frob}_{v})-T\cdot\text{Id})=P_{v}(T)\in E[T].$
The compatible system (3) is said to be semisimple if $\rho_{\lambda}$ is semisimple for all $\lambda$.
2.3. Local algebraicity
Define $\mathbf{T}_{K}:=\mathrm{Res}_{K/\mathbb{Q}}\mathbb{G}_{m,K}$ to be the $[K:\mathbb{Q}]$-dimensional $\mathbb{Q}$-torus given by restriction of scalars of $\mathbb{G}_{m,K}:=\mathrm{Spec}K[x,x^{-1}]$. For any commutative $\mathbb{Q}$-algebra $A$, its group of $A$-points is
$\mathbf{T}_{K}(A)=(K\otimes_{\mathbb{Q}}A)^{\times}.$
Let $\mathbb{A}_{K}^{\times}$ be the group of idèles of $K$ and $i_{\ell}$ be the natural composition
$i_{\ell}:(K\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell})^{\times}=\prod_{v|\ell}K_{v}^{\times}\to\mathbb{A}_{K}^{\times}/K^{\times}\stackrel{{\scriptstyle\mathrm{Art}_{K}}}{{\longrightarrow}}\operatorname{Gal}_{K}^{\text{ab}},$
where $\mathrm{Art}_{K}$ is the Artin reciprocity map.
An abelian semisimple $\ell$-adic representation
$\phi_{\ell}:\operatorname{Gal}_{K}^{\text{ab}}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$
is said to be locally algebraic [Se98, Chap. III] if there exists an $\overline{\mathbb{Q}}_{\ell}$-algebraic morphism
$r_{\ell}:\mathbf{T}_{K}\times_{\mathbb{Q}}\overline{\mathbb{Q}}_{\ell}\to\mathrm{GL}_{n,\overline{\mathbb{Q}}_{\ell}}$
such that for all $x$ close enough to $1$ in the $\ell$-adic Lie group $(K\otimes_{\mathbb{Q}}\mathbb{Q}_{\ell})^{\times}=\mathbf{T}_{K}(\mathbb{Q}_{\ell})$:
$\phi_{\ell}\circ i_{\ell}(x)=r_{\ell}(x^{-1}).$
Remark 2.1.
Local algebraicity of $\phi_{\ell}$ depends solely on the local representations
$\phi_{\ell}|_{\operatorname{Gal}_{K_{v}}}\hskip 14.45377pt\text{for all}~{}v\in S_{\ell}.$
Moreover, $\phi_{\ell}$ is locally algebraic if and only if these local representations are Hodge-Tate, see [Se98, Appendix of Chap. III].
2.4. Equivalence of three notions
The following three notions of an abelian semisimple
$\phi_{\ell}:\operatorname{Gal}_{K}^{\text{ab}}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$
are equivalent.
(Loc-alg): $\phi_{\ell}$ is locally algebraic.
($E$-SCS): $\phi_{\ell}\in\{\phi_{\lambda}\}_{\lambda}$ a semisimple abelian Serre compatible system of $K$ defined over some number field $E$ unramified outside some finite $S\subset\Sigma_{K}$.
($E$-rat): $\phi_{\ell}$ is $E$-rational for some number field $E\subset\overline{\mathbb{Q}}_{\ell}$.
For direction (Loc-alg) $\Rightarrow$ ($E$-SCS), it is shown in [Se98, Chap. III.2.3, proof of Theorem 2] that $\phi_{\ell}$ admits a factorization
(4) $\phi_{\ell}:\operatorname{Gal}_{K}^{\text{ab}}\stackrel{{\scriptstyle\epsilon_{\ell}}}{{\longrightarrow}}\mathbf{S}_{\mathfrak{m}}(\overline{\mathbb{Q}}_{\ell})\stackrel{{\scriptstyle\phi_{0}}}{{\longrightarrow}}\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell}),$
where
  • $\mathbf{S}_{\mathfrak{m}}$ is the Serre group (for some modulus $\mathfrak{m}$ of $K$) [^1] The identity component $\mathbf{T}_{\mathfrak{m}}$ of $\mathbf{S}_{\mathfrak{m}}$ is a quotient of $\mathbf{T}_{K}$ and the component group $\mathbf{S}_{\mathfrak{m}}/\mathbf{T}_{\mathfrak{m}}$ is isomorphic to the Ray class group of $\mathfrak{m}$., which is a diagonalizable group defined over $\mathbb{Q}$;
  • $\epsilon_{\ell}:\operatorname{Gal}_{K}^{\text{ab}}\to\mathbf{S}_{\mathfrak{m}}(\mathbb{Q}_{\ell})$ is the canonical morphism constructed (for every $\ell$) in [Se98, Chap. II.2.3];
  • $\phi_{0}$ is an $\overline{\mathbb{Q}}_{\ell}$-algebraic morphism, which we assume to be a direct sum of characters.
Since $\mathbf{S}_{\mathfrak{m}}$ is defined over $\mathbb{Q}$, we may assume $\phi_{0}$ to be defined over some number field $E$. Then the family
(5) $\{\phi_{\lambda}:\operatorname{Gal}_{K}^{\text{ab}}\stackrel{{\scriptstyle\epsilon_{\ell(\lambda)}}}{{\longrightarrow}}\mathbf{S}_{\mathfrak{m}}(E_{\lambda})\stackrel{{\scriptstyle\phi_{0}\times_{E}E_{\lambda}}}{{\longrightarrow}}\mathrm{GL}_{n}(E_{\lambda})\subset\mathrm{GL}_{n}(\overline{E}_{\lambda}):~{}\lambda\in\Sigma_{E}\}$
is the SCS we want.
The direction ($E$-SCS) $\Rightarrow$ ($E$-rat) is obvious. Finally for ($E$-rat) $\Rightarrow$ (Loc-alg), according to the strategy of Serre in [Se98, Chap. III.3] there are two steps: the first one is to show that $\phi_{\ell}$ is almost locally algebraic (i.e., $\phi_{\ell}^{N}$ is locally algebraic for some $N\in\mathbb{N}$) and the second one is to show that almost local algebraicity implies local algebraicity for an $E$-rational $\phi_{\ell}$ (see Proposition 2.12). The first step is a direct consequence of Theorem 2.2 below which is proven (see e.g., [He82, Theorem 7]) by a remarkable theorem of Waldschmidt on transcendental numbers [Wa81].
Theorem 2.2 (Waldschmidt).
Suppose $\tau_{\ell}:\operatorname{Gal}_{K}\to\overline{\mathbb{Q}}_{\ell}^{\times}$ is an one-dimensional $\ell$-adic representation unramified outside a finite set $S\subset\Sigma_{K}$ such that $\tau_{\ell}(\operatorname{Frob}_{v})\in\overline{\mathbb{Q}}$ for all $v\in\Sigma_{K}\setminus S$. Then there exists $N\in\mathbb{N}$ such that $\tau_{\ell}^{N}$ is locally algebraic.
2.5. Formal character
Let $F$ be an algebraically closed field and $\mathbf{G}\subset\mathrm{GL}_{n,F}$ a reductive subgroup. The formal character of $\mathbf{G}$ is defined as a subtorus
(6) $\mathbf{T}\subset\mathrm{GL}_{n,F}$
up to $\mathrm{GL}_{n}(F)$-conjugation such that $g\mathbf{T}g^{-1}$ is a maximal torus of $\mathbf{G}$ for some $g\in\mathrm{GL}_{n}(F)$.
After some conjugation, (6) can be diagonalized and thus admits some $\mathbb{Z}$-forms:
(7) $\mathbf{T}_{\mathbb{Z}}\subset\mathbb{G}_{m,\mathbb{Z}}^{n}.$
If $\mathbf{G}^{\prime}\subset\mathrm{GL}_{n,F^{\prime}}$ is another reductive subgroup ($F^{\prime}$ some algebraically closed field), we say that the formal characters of $\mathbf{G}$ and $\mathbf{G}^{\prime}$ are the same if they admit the same $\mathbb{Z}$-form (7) after some diagonalizations.
2.6. Weak abelian direct summands
Throughout this subsection, we denote by $F$ the algebraically closed field $\overline{\mathbb{Q}}_{\ell}$ for simplicity. Let $\rho:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(F)$ be a semisimple $\ell$-adic representation that is unramified almost everywhere. We investigate a notion weaker than (abelian) subrepresentation of $\rho$, called weak (abelian) direct summand. When the algebraic monodromy group $\mathbf{G}_{\rho}$ satisfies some hypotheses (e.g., $\mathbf{G}_{\rho}$ is connected), weak abelian direct summands of $\rho$ are completely characterized and the weak abelian part (i.e., the maximal weak abelian direct summand) of $\rho$ is well-defined (see Proposition 2.9).
2.6.1. Weak divisibility
Definition 2.3.
Let $\rho:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(F)$ and $\psi:\operatorname{Gal}_{K}\to\mathrm{GL}_{m}(F)$ be semisimple $\ell$-adic representations that are unramified almost everywhere. Denote by $S_{\rho},S_{\psi}\subset\Sigma_{K}$ the set of ramified places of $\rho$ and $\chi$, respectively. We say that $\psi$ is a weak direct summand of $\rho$ (or $\psi$ weakly divides $\rho$) if the set
(8) $S_{\psi\mid\rho}:=\{v\in\Sigma_{K}\setminus(S_{\rho}\cup S_{\psi}):\det(\psi(\operatorname{Frob}_{v})-T\cdot\text{Id})~{}\mathrm{\,divides~{}\,}\det(\rho(\operatorname{Frob}_{v})-T\cdot\text{Id})\}$
is of Dirichlet density one. If $\psi$ is abelian and weakly divides $\rho$, we say that $\psi$ is a weak abelian direct summand of $\rho$.
Similarly, one can define weak divisibility for a pair of semisimple Serre compatible systems of $K$ defined over $E$. Note that if $\psi$ is a subrepresentation of $\rho$ then $\psi$ weakly divides $\rho$.
Proposition 2.4.
If $\psi$ weakly divides $\rho$, then $S_{\psi\mid\rho}=\Sigma_{K}\setminus(S_{\rho}\cup S_{\psi})$.
Proof.
Let $n$ and $m$ be the degrees of $\rho$ and $\psi$. For $\sigma\in\operatorname{Gal}_{K}$, denote by $\operatorname{Syl}_{\lambda,\sigma}$ the Sylvester matrix in $M_{(n+m)\times(n+m)}(E_{\lambda})$ for the polynomials $r_{\lambda,\sigma}(T):=\det(\rho(\sigma)-T\cdot\text{Id})$ and $s_{\lambda,\sigma}(T):=\det(\psi(\sigma)-T\cdot\text{Id})$ in $F[T]$ of degrees $n$ and $m$, respectively. By elementary linear algebra one has the following well-known result
(9) $\operatorname{rank}\operatorname{Syl}_{\lambda,\sigma}=n+m-\deg\gcd(r_{\lambda,\sigma},s_{\lambda,\sigma}).$
Fix $r$ row and column labels in $\{1,\ldots,n+m\}$, i.e., $r$ element subsets $I,J\subset\{1,\ldots,n+m\}$, and consider the determinant of the $r\times r$ minor of $\operatorname{Syl}_{\lambda,\sigma}$ spanned by the chosen rows and columns; call it $d_{\lambda,I,J}(\sigma)$. Then for each $\lambda,I,J$ the map $d_{\lambda,I,J}:G_{K}\to E_{\lambda}$ that sends $\sigma\mapsto d_{\lambda,I,J}(\sigma)$ is continuous, and by our hypothesis, whenever $|I|=|J|>n$, the map vanishes on the elements $\operatorname{Frob}_{v}$, for $v$ in a density one subset of places of $K$. Hence by the Chebotarev density theorem the map $d_{\lambda,I,J}$ is identically zero for any $|I|=|J|>n$. It follows that $s_{\lambda,\sigma}$ divides $r_{\lambda,\sigma}$ for all $\sigma\in\operatorname{Gal}_{K}$, and choosing $\sigma=\operatorname{Frob}_{v}$ for $v\in\Sigma_{K}\setminus(S_{\rho}\cup S_{\psi})$ and $\lambda$ a place of $E$ with residue characteristic different from $v$, the result follows. ∎
Remark 2.5.
The notion of weak divisibility from Definition 2.3 has some pitfalls that we’d like to point out: It is easy to see that if $\psi$ weakly divides $\rho$ then the representation $\psi$ need not be a direct summand of $\rho$; for instance, the trivial $1$-dimensional representation is a weak direct summand of any $3$-dimensional $\ell$-adic representation with image in $\mathrm{SO}_{3}$. But in fact one can have $\rho$ and $\psi$ abelian and semisimple with $\psi$ weakly dividing $\rho$, and yet $\psi$ not being a direct summand of $\rho$. Let $L/K$ be a finite Galois extension with $\operatorname{Gal}(L/K)\cong\mathbb{Z}/2\times\mathbb{Z}/2$, let $\chi_{i}$, $i=0,1,2,3$, denote the characters $\operatorname{Gal}(L/K)\to\{\pm 1\}$. Then any $\chi_{i}$ is a weak direct summand of the sum of the other $3$ characters.
In the abelian case, this type of problem can arise whenever there are characters $\chi$ and $\chi^{\prime}$ that are direct summands of $\psi$ and $\rho$ such that $\chi^{\prime}\otimes\chi^{-1}$ has finite non-trivial order. Ruling out this case, below in Proposition 2.9 and Corollary 2.10 we shall analyze in detail when an abelian $\psi$ weakly divides $\rho$.
2.6.2. Some lemmas
For the next results we denote by $\mathbb{G}_{m}$ the center of $\mathrm{GL}_{n,F}$, and by $\mathbbold{1}:\operatorname{Gal}_{K}\to\mathrm{GL}_{1}(F)$ the trivial character. For a reductive group $\mathbf{G}$ defined over $F$, denote by $\mathbf{G}^{\operatorname{der}}$ the derived group of the identity component $\mathbf{G}^{\circ}$.
Lemma 2.6.
Let $\mathbf{G}\subset\mathrm{GL}_{n,F}$ be a closed subgroup that acts irreducible on $F^{n}$ and suppose that $\mathbf{G}/Z(\mathbf{G})$ is connected, where $Z(\mathbf{G})\subset\mathbf{G}$ is the center. Let $n_{0}$ be the multiplicity of the weight zero in the formal character of $\mathbf{G}^{\operatorname{der}}$ on $F^{n}$. Then the following hold:
  1. (i)
    The center $Z(\mathbf{G})$ lies in $\mathbb{G}_{m}$, and so is equal to $\mathbf{G}\cap\mathbb{G}_{m}$, an element $\zeta\cdot 1$ in the center acts by multiplication with $\zeta$ on $F^{n}$. Moreover, if $Z(\mathbf{G})$ is finite then $\mathbf{G}^{\circ}$ is semisimple. In all cases we have $\mathbf{G}=(\mathbf{G}\cap\mathbb{G}_{m})\mathbf{G}^{\operatorname{der}}$.
  2. (ii)
    Suppose $n_{0}>0$. Then we have: $\mathbf{G}^{\operatorname{der}}\cap\mathbb{G}_{m}=\{1\}$ and a direct product
    $\mathbf{G}=\mathbf{G}^{\operatorname{der}}\times(\mathbf{G}\cap\mathbb{G}_{m}).$
    Moreover denoting by $\nu_{\mathbf{G}}:\mathbf{G}\to\mathbf{G}\cap\mathbb{G}_{m}$ the projection onto the second factor, one has $\ker(\nu_{\mathbf{G}})=\mathbf{G}^{\operatorname{der}}$ and $\det=(\nu_{\mathbf{G}})^{\otimes n}$ with $\det$ the map $\mathbf{G}\subset\mathrm{GL}_{n}\stackrel{{\scriptstyle\det}}{{\to}}\mathbb{G}_{m}$.
Proof.
The first sentence of (i) follows from the irreducibility of the action of $\mathbf{G}$ on $F^{n}$. The connectedness of $\mathbf{G}/Z(\mathbf{G})$ implies that $\mathbf{G}=Z(\mathbf{G})\mathbf{G}^{\circ}$. It follows that $\mathbf{G}^{\circ}$ acts irreducibly and is thus reductive connected. In particular $Z(\mathbf{G}^{\circ})$ lies in $\mathbb{G}_{m}$ and one has $Z(\mathbf{G}^{\circ})\mathbf{G}^{\operatorname{der}}=\mathbf{G}^{\circ}$. Therefore, if $\mathbf{G}\cap\mathbb{G}_{m}$ is finite, so must be $\mathbf{G}^{\circ}\cap\mathbb{G}_{m}=Z(\mathbf{G}^{\circ})$ and then $\mathbf{G}^{\circ}$ is semisimple and $\mathbf{G}=(\mathbf{G}^{\circ}\cap\mathbb{G}_{m})\mathbf{G}^{\operatorname{der}}$. Otherwise, $\mathbf{G}$ contains $\mathbb{G}_{m}$ and we have $\mathbf{G}=\mathbf{G}^{\circ}=\mathbb{G}_{m}\mathbf{G}^{\operatorname{der}}$.
For the proof of (ii) assume $n_{0}>0$. Since the intersection $\mathbf{G}^{\operatorname{der}}\cap\mathbb{G}_{m}$ cannot contain any non-trivial scalar matrices, it follows that $\mathbf{G}^{\operatorname{der}}\cap\mathbb{G}_{m}=\{1\}$. By (i) the center $Z(\mathbf{G})$ together with $\mathbf{G}^{\operatorname{der}}$ generate $\mathbf{G}$, and they are both normal subgroups of $\mathbf{G}$. As we have shown, they intersect trivially, and now part (ii) follows from elementary group theory. The last assertions, on $\nu_{\mathbf{G}}$, are then straightforward. ∎
Lemma 2.7.
Let $\rho\colon\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(F)$ and $\chi:\operatorname{Gal}_{K}\to\mathrm{GL}_{1}(F)$ be semisimple $\ell$-adic representations. Denote by $\mathbf{G}_{\rho}$ and $\mathbf{G}_{\rho\otimes\chi}$ respectively the algebraic monodromy groups of $\rho$ and $\rho\otimes\chi$. Then the following assertions hold.
  1. (i)
    The quotient $\mathbf{G}_{\rho}/Z(\mathbf{G}_{\rho})$ is connected if and only if $\mathbf{G}_{\rho\otimes\chi}/Z(\mathbf{G}_{\rho\otimes\chi})$ is connected.
  2. (ii)
    The multiplicities of the weight zero in the formal characters of $\mathbf{G}_{\rho}^{\operatorname{der}}$ and $\mathbf{G}_{\rho\otimes\chi}^{\operatorname{der}}$ on $F^{n}$ are equal.
Proof.
The lemma follows from the facts that $\mathbf{G}_{\rho}/Z(\mathbf{G}_{\rho})\simeq\mathbf{G}_{\rho\otimes\chi}/Z(\mathbf{G}_{\rho\otimes\chi})$ and the actions of $\mathbf{G}_{\rho}^{\operatorname{der}}$ and $\mathbf{G}_{\rho\otimes\chi}^{\operatorname{der}}$ on $F^{n}$ can be identified. ∎
To proceed, we recall the following result of Steinberg, e.g. [GW09, Thm. 11.4.18], that implicitly made its appearance already in the proof of Proposition 2.11 when analyzing the set $\mathbf{Y}_{\ell}$. If $\mathbf{G}$ is connected reductive group over $F$, then the conjugation-invariant
(10) $\mathbf{G}^{\mathrm{reg.ss.}}:=\{g\in\mathbf{G}\mid g\hbox{ is regular semisimple}\}\hbox{ is a dense open subset of }\mathbf{G}.$
For a subset $T\subset\Sigma_{K}$, we write $\delta(T)$ for its (Dirichlet) density if it exists.
Lemma 2.8.
Let $\rho\colon\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(F)$ and $\chi:\operatorname{Gal}_{K}\to\mathrm{GL}_{1}(F)$ be $\ell$-adic representations that are both unramified outside a finite subset $S\subset\Sigma_{K}$. Denote by $\mathbf{G}_{\rho}\subset\mathrm{GL}_{n,F}$ the algebraic monodromy group of $\rho$ and by $n_{0}$ the multiplicity of the weight zero in the formal character of $\mathbf{G}_{\rho}^{\operatorname{der}}$ on $F^{n}$. Suppose that $\rho$ is irreducible and $\mathbf{G}_{\rho}/Z(\mathbf{G}_{\rho})$ is connected. If $n_{0}>0$, define the character $\nu_{\mathbf{G}_{\rho}}:\operatorname{Gal}_{K}\to(\mathbf{G}_{\rho}\cap\mathbb{G}_{m})(F^{\times})\subset F^{\times}$. Then the following assertions hold.
  1. (i)
    If $n_{0}=0$, then $\delta(S_{\chi|\rho})=0$.
  2. (ii)
    If $n_{0}>0$ and the order of $\det(\rho\otimes\chi^{-1})$ is infinite, then $\mathbb{G}_{m}\subset\mathbf{G}_{\rho\otimes\chi^{-1}}$ and $\delta(S_{\chi|\rho})=0$.
  3. (iii)
    If $n_{0}>0$ and the order of $\det(\rho\otimes\chi^{-1})$ is finite, then $Z(\mathbf{G}_{\rho\otimes\chi^{-1}})=\mathbf{G}_{\rho\otimes\chi^{-1}}\cap\mathbb{G}_{m}$ is finite, the symmetric difference of $S_{\chi|\rho}$ and $S_{\mathbbold{1}|\nu_{\mathbf{G}_{\rho\otimes\chi^{-1}}}}$ is of density zero and
    $\delta(S_{\chi|\rho})=\delta(S_{\mathbbold{1}|\nu_{\mathbf{G}_{\rho\otimes\chi^{-1}}}})=\frac{1}{|Z(\mathbf{G}_{\rho\otimes\chi^{-1}})|}.$
  4. (iv)
    The character $\chi$ weakly divides $\rho$ if and only if $n_{0}>0$ and $\chi=\nu_{\mathbf{G}_{\rho}}$.
  5. (v)
    If $\chi$ weakly divides $\rho$, then $\delta(S_{\chi^{\oplus n_{0}}|\rho})=1$ and $\delta(S_{\chi^{\oplus(n_{0}+1)}|\rho})=0$.
Proof.
Since the symmetric difference of $S_{\chi|\rho}$ and $S_{\mathbbold{1}|\rho\otimes\chi^{-1}}$ is finite, it follows that $\delta(S_{\chi|\rho})=\delta(S_{\mathbbold{1}|\rho\otimes\chi^{-1}})$. As all our assertions concern $\rho\otimes\chi^{-1}$, we may assume $\chi=\mathbbold{1}$ by Lemma 2.7. Note also that $\mathbf{G}=\mathbf{G}_{\rho}$ satisfies the hypotheses of Lemma 2.6, so that the lemma applies.
Suppose first that $\mathbf{G}_{\rho}$ contains $\mathbb{G}_{m}$, which happens exactly when $\det(\rho)$ has infinite order. Then $\mathbf{G}_{\rho}$ is connected reductive and has center $\mathbb{G}_{m}$. The conjugation-invariant subvariety
$Y_{1}:=\{g\in\mathbf{G}:1\hbox{ is an eigenvalue of }g\}\subset\mathbf{G}_{\rho}$
is closed and no $g\in Y_{1}$ is regular semisimple because $\mathbb{G}_{m}$ does not lie in the closure of $g^{\mathbb{Z}}$. Hence, (10) implies that $Y_{1}$ is properly contained in $\mathbf{G}_{\rho}$. Because $\mathbf{G}_{\rho}$ is connected, it follows that $S_{1}:=\{v\in\Sigma_{K}\backslash S:\rho(\operatorname{Frob}_{v})\in Y_{1}\}$ has density zero, and hence $\delta(S_{\mathbbold{1}|\rho})=0$. Thus, we obtain (i) in case $\mathbb{G}_{m}\subset\mathbf{G}_{\rho}$ and also (ii).
Suppose now that $\mathbf{G}_{\rho}$ does not contain $\mathbb{G}_{m}$. Then $\mathbf{G}_{\rho}^{\circ}=\mathbf{G}_{\rho}^{\operatorname{der}}$ and $\mathbf{G}_{\rho}\cap\mathbb{G}_{m}=Z(\mathbf{G}_{\rho})$ is a finite subgroup $\mu_{N}=\{\zeta\in F:\zeta^{N}=1\}$ of $\mathbb{G}_{m}$, where $N=|Z(\mathbf{G}_{\rho})|\in\mathbb{N}$. Because $\mathbf{G}_{\rho}=Z(\mathbf{G}_{\rho})\mathbf{G}_{\rho}^{\operatorname{der}}$, any element in $\mathbf{G}_{\rho}$ can be written as $\zeta g$ with $\zeta\in\mu_{N}$ and $g\in\mathbf{G}_{\rho}^{\operatorname{der}}$. Now from (10) it follows that
$\{g\in\mathbf{G}_{\rho}^{\operatorname{der}}:\hbox{ some eigenvalue of }g~{}\hbox{belongs to }\mu_{N}\backslash\{1\}\}$
is a conjugation-invariant proper closed subvariety of $\mathbf{G}_{\rho}^{\operatorname{der}}$, and therefore
(11) $Y_{2}:=\{g\in\mathbf{G}_{\rho}\setminus\mathbf{G}_{\rho}^{\circ}:1\hbox{ is aneigenvalue of }g\}$
is a conjugation-invariant closed nowhere dense subvariety of $\mathbf{G}_{\rho}\setminus\mathbf{G}^{\circ}_{\rho}$. Hence, the Chebotarev density theorem implies that the set $S_{2}:=\{v\in\Sigma_{K}\backslash S:\rho(\operatorname{Frob}_{v})\in Y_{2}\}$ has density zero. By the definition of $n_{0}$, the conjugation-invariant
(12) $Y_{3}:=\{g\in\mathbf{G}_{\rho}^{\circ}:1\hbox{ is an eigenvalue of }g\}\subset\mathbf{G}_{\rho}^{\circ}$
is all of $\mathbf{G}_{\rho}^{\circ}$ if $n_{0}>0$ and is proper closed subvariety if $n_{0}=0$. Hence, the Chebotarev density theorem implies that the set $S_{3}:=\{v\in\Sigma_{K}\backslash S:\rho(\operatorname{Frob}_{v})\in Y_{3}\}$ has density $1/[\mathbf{G}_{\rho}:\mathbf{G}^{\circ}_{\rho}]$ if $n_{0}>0$ and density zero if $n_{0}=0$. In case $n_{0}=0$, we obtain that $\delta(S_{\mathbbold{1}|\rho})=\delta(S_{2})+\delta(S_{3})=0$, and this completes (i). In case $n_{0}>0$, Lemma 2.6 yields
$[\mathbf{G}_{\rho}:\mathbf{G}^{0}_{\rho}]=|\mathbb{G}_{m}\cap\mathbf{G}_{\rho}|=|Z(\mathbf{G}_{\rho})|=N.$
By the above results on $\delta(S_{2})$ and $\delta(S_{3})$, we deduce (iii), and in particular
$\delta(S_{\mathbbold{1}|\rho})=\delta(S_{2})+\delta(S_{3})=\frac{1}{N}=\delta(S_{\mathbbold{1}|\nu_{\mathbf{G}_{\rho}}}).$
[…]
(13) $Y_{4}:=\{g\in\mathbf{G}_{\rho}:1\hbox{ is an eigenvalue of }g\hbox{ of multiplicity at least $n_{0}+1$}\}$
contains no regular semisimple elements, and by (10) it is nowhere dense in $\mathbf{G}_{\rho}$. By the Chebotarev density theorem, we deduce that $\{v\in\Sigma_{K}\backslash S:\rho(\operatorname{Frob}_{v})\in Y_{4}\}=S_{\mathbbold{1}^{\oplus(n_{0}+1)}|\rho}$ has density zero. ∎
2.6.3. Characterization of weak abelian direct summands and the weak abelian part
Proposition 2.9.
Let $\rho\colon\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(F)$ be a semisimple $\ell$-adic representation unramified outside a finite subset $S\subset\Sigma_{K}$ that can be written as $\rho=\bigoplus_{j\in J}\rho_{j}$ with absolutely irreducible representations $\rho_{j}:\operatorname{Gal}_{K}\to\mathrm{GL}_{r_{j}}(F)$ and suppose that for each $j\in J$ the group $\mathbf{G}_{\rho_{j}}/Z(\mathbf{G}_{\rho_{j}})$ is connected, where $\mathbf{G}_{\rho_{j}}$ is the algebraic monodromy group of $\rho_{j}$. Denote by $n_{j,0}$ the multiplicity of the weight zero in the formal character of $\mathbf{G}_{\rho_{j}}^{\operatorname{der}}$ on $F^{r_{j}}$, let $J_{0}:=\{j\in J:n_{j,0}>0\}$ and for $j\in J_{0}$ define $\xi_{j}:=\nu_{\mathbf{G}_{\rho_{j}}}$ the Galois character in Lemma 2.8. Suppose for all $j,j^{\prime}\in J_{0}$ that $\xi_{j^{\prime}}\xi_{j}^{-1}$ is either trivial or has infinite order. Define a partition $J_{0}=\bigsqcup_{a\in A}J_{a}$ be requiring $j,j^{\prime}$ to lie in the same class if and only if $\xi_{j}=\xi_{j^{\prime}}$, and write $\xi_{a}$ for $\xi_{j}$ if $j\in J_{a}$.
Let also $\psi=\bigoplus_{i\in I}\psi_{i}^{\oplus e_{i}}$ be an abelian $\ell$-adic representation with continuous characters $\psi_{i}:\operatorname{Gal}_{K}\to F^{\times}$ unramified outside $S$, that are pairwise distinct and with $e_{i}\in\mathbb{N}$. Then $\psi$ weakly divides $\rho$ if and only if the following conditions hold:
  1. (a)
    for each $i\in I$ there is an (unique) $a_{i}\in A$ such that $\psi_{i}=\xi_{a_{i}}$, and
  2. (b)
    for all $i\in I$ one has $e_{i}\leq\sum_{j\in J_{a_{i}}}n_{j_{i},0}$.
In particular, there is a unique degree $n_{0}:=\sum_{j\in J}n_{j,0}$ weak abelian direct summand $\rho^{\text{wab}}$ of $\rho$ such that any weak abelian direct summand $\psi$ of $\rho$ is a subrepresentation of $\rho^{\text{wab}}$. We call $\rho^{\text{wab}}$ the weak abelian part of $\rho$.
Moreover if $\psi$ weakly divides $\rho$ and $V_{\psi}$ is the representation space of $\psi$, then there is an algebraic representation $r_{\psi}:\mathbf{G}_{\rho}\to\mathrm{GL}_{V_{\psi}}$ such that $\psi$ is the composition
(14) $\operatorname{Gal}_{K}\stackrel{{\scriptstyle\rho}}{{\rightarrow}}\rho(\operatorname{Gal}_{K})\hookrightarrow\mathbf{G}_{\rho}(F)\stackrel{{\scriptstyle r_{\psi}}}{{\rightarrow}}\mathrm{GL}_{V_{\psi}}(F)=\mathrm{GL}(V_{\psi}).$
In particular, $r_{\psi}(\mathbf{G}_{\rho})=\mathbf{G}_{\psi}$ is the algebraic monodromy group of $\psi$.
Proof.
Suppose first that (a) and (b) hold. Then from Lemma 2.8, it follows that $\psi_{i}^{\oplus e_{i}}$ weakly divides $\bigoplus_{j\in J_{a_{i}}}\rho_{j}$. Summing over all $i$ gives that $\psi$ weakly divides $\rho$.
Suppose conversely that $\psi$ weakly divides $\rho$. Then in particular for each $i$ we have $\delta(S_{\psi_{i}|\rho})=1$. If $j\in J\setminus J_{0}$, then $\delta(S_{\psi_{i}|\rho_{j}})$ is zero by Lemma 2.8(i), because $n_{j,0}=0$. If $j\in J_{0}$, then $\xi_{j}$ is defined. By our hypothesis on the $\xi_{j}$, the order of $\xi_{a}\otimes\psi_{i}^{-1}$ can be finite for at most one $a\in A$. If the order is infinite, by Lemma 2.8(ii) we have $\delta(S_{\psi_{i}|\rho_{j}})=0$. Therefore there must be (exactly) one $a\in A$ such that $\delta(S_{\psi_{i}|\bigoplus_{j\in J_{a}}\rho_{j}})=1$. From Lemma 2.8(iii) we deduce that the symmetric difference of $S_{\psi_{i}|\rho_{j}}$ and $S_{\psi_{i}|\xi_{a}}$ is of density zero and that $\delta(S_{\psi_{i}|\xi_{a}})=\frac{1}{\mathrm{ord}(\psi_{i}\xi_{a}^{-1})}$ for $j\in J_{a}$. It follows that
$\frac{1}{\mathrm{ord}(\psi_{i}\xi_{a}^{-1})}=\delta(S_{\psi_{i}|\xi_{a}})=\delta(\bigcup_{j\in J_{a}}S_{\psi_{i}|\rho_{j}})=\delta(S_{\psi_{i}|\bigoplus_{j\in J_{a}}\rho_{j}})=1$
and we must have $\psi_{i}=\xi_{a}$ for an (unique) $a\in A$, which we now denote by $a_{i}$.
It remains to prove (b). From the above we have $\delta(S_{\psi_{i}|\bigoplus_{j\in J_{a_{i}}}\rho_{j}})=1$. We know that
$S_{\psi_{i}^{\oplus e_{i}}|\bigoplus_{j\in J_{a_{i}}}\rho_{j}}=\bigcup_{(d_{j})_{j\in J_{a_{i}}}}\bigcap_{j\in J_{a_{i}}}S_{\psi_{i}^{\oplus d_{j}}|\rho_{j}}$
where the union is over all $J_{a_{i}}$-indexed tuples $(d_{j})$ of elements in $\mathbb{Z}_{\geq 0}$ such that $\sum_{j\in J_{a_{i}}}d_{j}=e_{i}$. If $d_{j}>n_{j,0}$, then $\delta(S_{\psi_{i}^{\oplus d_{j}}|\rho_{j}})=0$ by Lemma 2.8(v); otherwise the density is one. Hence there must be a tuple $(d_{j})$ such that $\delta(S_{\psi_{i}^{\oplus d_{j}}|\rho_{j}})=1$ for all $j\in J_{a_{i}}$. But then $d_{j}\leq n_{j,0}$, and so
(15) $e_{i}=\sum_{j\in J_{a_{i}}}d_{j}\leq\sum_{j\in J_{a_{i}}}n_{j,0},$
as was to be shown. When the equality of (15) holds for all $i$, one obtains the weak abelian part $\rho^{\text{wab}}=\psi$ of $\rho$ with the desired properties.
Finally note that if $\psi$ weakly divides $\rho$, we obtain $r_{\psi}$ and (14) by condition (a) and (the direct product in) Lemma 2.6(ii). Since $\rho(\operatorname{Gal}_{K})$ is Zariski dense in $\mathbf{G}_{\rho}$, (14) implies that $\psi(\operatorname{Gal}_{K})$ is Zariski dense in $r_{\psi}(\mathbf{G}_{\rho})$ and thus $\mathbf{G}_{\psi}=r_{\psi}(\mathbf{G}_{\rho})$. ∎
Corollary 2.10.
Let $\rho\colon\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(F)$ and $\psi:\operatorname{Gal}_{K}\to\mathrm{GL}_{m}(F)$ be semisimple $\ell$-adic representations unramified outside a finite subset $S\subset\Sigma_{K}$, and suppose that $G_{\rho}$ is connected. Then the following assertions hold.
  1. (i)
    The hypotheses of Proposition 2.9 on $\rho$ are satisfied under the connectedness of $\mathbf{G}_{\rho}$.
  2. (ii)
    If $\psi$ is abelian and weakly divides $\rho$, then the algebraic monodromy $\mathbf{G}_{\psi}$ of $\psi$ is connected.
  3. (iii)
    If $\rho$ and $\psi$ are abelian, then $\psi$ weakly divides $\rho$ if and only if $\psi$ is a direct summand of $\rho$.
Proof.
Since $\mathbf{G}_{\rho}$ is connected, it is obvious that $\mathbf{G}_{\rho_{j}}/Z(\mathbf{G}_{\rho_{j}})$ is connected for all $j\in J$. Suppose $\xi_{j}$ and $\xi_{j^{\prime}}$ are two Galois characters in Proposition 2.9 where $j,j^{\prime}\in J_{0}$, and let $\mathbf{G}_{jj^{\prime}}\subset\mathbb{G}_{m}$ be the algebraic monodromy of $\xi_{j}\xi_{j^{\prime}}^{-1}$. By Lemma 2.6(ii), there is an algebraic representation $r_{jj^{\prime}}:\mathbf{G}_{\rho}\to\mathbb{G}_{m}$ such that $\xi_{j}\xi_{j^{\prime}}^{-1}$ is the composition
(16) $\operatorname{Gal}_{K}\stackrel{{\scriptstyle\rho}}{{\rightarrow}}\rho(\operatorname{Gal}_{K})\hookrightarrow\mathbf{G}_{\rho}(F)\stackrel{{\scriptstyle r_{jj^{\prime}}}}{{\rightarrow}}\mathbb{G}_{m}(F)=F^{\times}.$
Since $\rho(\operatorname{Gal}_{K})$ is Zariski dense in $\mathbf{G}_{\rho}$, (16) implies that $\xi_{j}\xi_{j^{\prime}}^{-1}(\operatorname{Gal}_{K})$ is Zariski dense in $r_{jj^{\prime}}(\mathbf{G}_{\rho})$, which is $\mathbf{G}_{jj^{\prime}}$. It follows that $\mathbf{G}_{jj^{\prime}}$ is also connected and thus, $\xi_{j}\xi_{j^{\prime}}^{-1}$ is either trivial or infinite. We obtain (i).
By (i) and (the last assertion of) Proposition 2.9, we obtain a surjection $\mathbf{G}_{\rho}\to\mathbf{G}_{\psi}$ of algebraic groups. Since $\mathbf{G}_{\rho}$ is connected, $\mathbf{G}_{\psi}$ is also connected, which is (ii). Finally, (iii) follows directly from (i) and Proposition 2.9 since $\rho$ is equal to its weak abelian part $\rho^{\text{wab}}$. ∎
2.7. Local algebraicity of weak abelian direct summand
Let $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ be an $E$-rational semisimple $\ell$-adic representation for some number field $E\subset\overline{\mathbb{Q}}_{\ell}$ and let $\psi_{\ell}$ be a weak abelian direct summand of $\rho_{\ell}$. Since $\rho_{\ell}$ is semisimple, the algebraic monodromy group $\mathbf{G}_{\ell}$ of $\rho_{\ell}$ is a reductive subgroup of $\mathrm{GL}_{n,\overline{\mathbb{Q}}_{\ell}}$. We are now ready to prove that $\psi_{\ell}$ is locally algebraic (Theorem 1.2).
Proof of Theorem 1.2. For any finite field extension $L/K$, the restriction $\rho_{\ell}|_{\operatorname{Gal}_{L}}$ is also $E$-rational. Since the local algebraicity of $\psi_{\ell}$ follows from the local algebraicity of $\psi_{\ell}|_{\operatorname{Gal}_{L}}$, we may assume that the algebraic monodromy group $\mathbf{G}_{\ell}$ is connected by replacing $K$ by some finite extension. Since $\psi_{\ell}$ weakly divides $\rho_{\ell}$, there exists $N\in\mathbb{N}$ such that $\psi_{\ell}^{N}$ is locally algebraic by $E$-rationality of $\rho_{\ell}$ and Theorem 2.2. Then the implication (Loc-alg) $\Rightarrow$ ($E$-rat) for semisimple abelian $\ell$-adic representations implies the existence of a number field $E^{\prime}\subset\overline{\mathbb{Q}}_{\ell}$ such that $\psi_{\ell}^{N}$ is $E^{\prime}$-rational. By Propositions 2.11 and 2.12 below, $\psi_{\ell}$ is locally algebraic. ∎
Proposition 2.11.
Let $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ be a semisimple $E^{\prime}$-rational $\ell$-adic representation unramified outside a finite $S\subset\Sigma_{K}$ such that the algebraic monodromy group $\mathbf{G}_{\ell}$ is connected. If $\psi_{\ell}$ is a weak abelian direct summand of $\rho_{\ell}$ and there exists $N\in\mathbb{N}$ such that $\psi_{\ell}^{N}$ is $E^{\prime}$-rational, then there is a Dirichlet density one subset $\mathcal{L}_{K}\subset\Sigma_{K}\backslash S$ such that
(17) $\det(\psi_{\ell}(\operatorname{Frob}_{v})-T\cdot\mathrm{Id})\in E^{\prime}[T]\hskip 14.45377pt\text{for all}~{}v\in\mathcal{L}_{K}.$
Proof.
As $\psi_{\ell}$ is abelian, write
(18) $\psi_{\ell}=\chi_{1}\oplus\chi_{2}\oplus\cdots\oplus\chi_{m}$
as a direct sum of characters. Since $\psi_{\ell}^{N}$ is $E^{\prime}$-rational, for all $v$ outside some finite $S^{\prime}\subset\Sigma_{K}$ the polynomial
(19) $\prod_{i=1}^{m}(T^{N}-\chi_{i}(\operatorname{Frob}_{v})^{N})\in E^{\prime}[T].$
Let $V_{\rho_{\ell}}$ and $V_{\psi_{\ell}^{-1}}$ be the representation spaces of $\rho_{\ell}$ and $\psi_{\ell}^{-1}$ respectively. Denote by $i_{\ell}$ the faithful representation of $\mathbf{G}_{\ell}$ on $V_{\rho_{\ell}}$. Since $\mathbf{G}_{\ell}$ is connected, the hypotheses of Proposition 2.9 on $\rho_{\ell}$ are satisfied by Corollary 2.10(i). Thus, Proposition 2.9 asserts that the twisted $\ell$-adic representation $\rho_{\ell}\otimes\psi_{\ell}^{-1}$ is equal to the composition
$\operatorname{Gal}_{K}\stackrel{{\scriptstyle\rho_{\ell}}}{{\rightarrow}}\rho_{\ell}(\operatorname{Gal}_{K})\hookrightarrow\mathbf{G}_{\ell}(\overline{\mathbb{Q}}_{\ell})\stackrel{{\scriptstyle i_{\ell}\otimes r_{\psi_{\ell}}^{-1}}}{{\longrightarrow}}\mathrm{GL}_{V_{\rho_{\ell}}\otimes V_{\psi_{\ell}^{-1}}}(\overline{\mathbb{Q}}_{\ell})=\mathrm{GL}(V_{\rho_{\ell}}\otimes V_{\psi_{\ell}^{-1}}),$
where $r_{\psi_{\ell}}$ is as in (14). Hence, the algebraic monodromy group $\mathbf{H}_{\ell}$ of $\rho_{\ell}\otimes\psi_{\ell}^{-1}$ is also connected reductive. Let $\mu_{N}\subset\overline{\mathbb{Q}}_{\ell}^{\times}$ be the group of $N$th root of unity. Define the conjugation-invariant
$\mathbf{Y}_{\ell}:=\{g\in\mathbf{H}_{\ell}:~{}\text{some eigenvalue of }g~{}\text{belongs to~{}}\mu_{N}\backslash\{1\}\}.$
Then $\mathbf{Y}_{\ell}$ is a proper closed subvariety of $\mathbf{H}_{\ell}$ as it does not contain the identity. Pick a large enough intermediate local field $\mathbb{Q}_{\ell}\subset L_{\ell}\subset\overline{\mathbb{Q}}_{\ell}$ such that $\rho_{\ell}\otimes\psi_{\ell}^{-1}(\operatorname{Gal}_{K})\subset\mathrm{GL}_{mn}(L_{\ell})$ (so that $\mathbf{H}_{\ell}$ defined over $L_{\ell}$) and $\mathbf{Y}_{\ell}$ is defined over $L_{\ell}$. The (algebraic) Chebotarev density theorem ([Se81a],[Raj98, Theorem 3]) implies that
(20) $\mathcal{L}_{K}:=\{v\in\Sigma_{K}\backslash S^{\prime}:~{}\rho_{\ell}\otimes\psi_{\ell}^{-1}(\operatorname{Frob}_{v})\in\mathbf{H}_{\ell}(L_{\ell})\backslash\mathbf{Y}_{\ell}(L_{\ell})\}$
is of natural (hence Dirichlet) density one. Therefore, for any $v\in\mathcal{L}_{K}$ the polynomial
$\prod_{i=1}^{m}(T-\chi_{i}(\operatorname{Frob}_{v}))\in E^{\prime}[T]$
because it is the greatest common divisor of (19) and $\det(\rho_{\ell}(\operatorname{Frob}_{v})-T\cdot\text{Id})$ in $E^{\prime}[T]$. We are done. ∎
Below is a slight modification of [Se98, Chap. III.3.2, Proposition].
Proposition 2.12 (Serre).
Suppose $\phi_{\ell}:\operatorname{Gal}_{K}^{\mathrm{ab}}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ is an abelian semisimple $\ell$-adic representation unramified outside a finite $S\subset\Sigma_{K}$ such that $\phi_{\ell}^{N}$ is locally algebraic for some $N\in\mathbb{N}$. If there is a number field $E^{\prime}\subset\overline{\mathbb{Q}}_{\ell}$ and a Dirichlet density one subset $\mathcal{L}_{K}\subset\Sigma_{K}\backslash S$ such that
(21) $\det(\phi_{\ell}(\operatorname{Frob}_{v})-T\cdot\mathrm{Id})\in E^{\prime}[T]\hskip 14.45377pt\text{for all}~{}v\in\mathcal{L}_{K},$
then $\phi_{\ell}$ is locally algebraic and $E^{\prime}$-rational.
Proof.
[…]
$\phi_{\ell}:\operatorname{Gal}_{K}^{\mathrm{ab}}\stackrel{{\scriptstyle\epsilon_{\ell}}}{{\longrightarrow}}\mathbf{S}_{\mathfrak{m}}(\overline{\mathbb{Q}}_{\ell})\stackrel{{\scriptstyle\phi_{0}}}{{\longrightarrow}}\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell}).$
Since the trace $\mathrm{Tr}(\phi_{\ell}(\operatorname{Frob}_{v}))\in E^{\prime}$ for all $v\in\mathcal{L}_{K}$ by (21) and the subset
$\{\phi_{\ell}(\operatorname{Frob}_{v}):~{}v\in\mathcal{L}_{K}\}\subset\phi_{\ell}(\operatorname{Gal}_{K})$
is dense by $\mathcal{L}_{K}$ of Dirichlet density one, the morphism $\phi_{0}$ is defined over $E^{\prime}$ [Se98, Chap. II.2.4, Proposition 2 and Lemma]. Therefore, $\phi_{\ell}$ is $E^{\prime}$-rational. ∎
Corollary 2.13.
Let $\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})$ be a semisimple $E^{\prime}$-rational $\ell$-adic representation unramified outside a finite $S\subset\Sigma_{K}$ such that the algebraic monodromy group $\mathbf{G}_{\ell}$ is connected. If $\psi_{\ell}$ is a weak abelian direct summand of $\rho_{\ell}$ and there exists a finite field extension $L/K$ such that $\psi_{\ell}|_{\operatorname{Gal}_{L}}$ is $E^{\prime}$-rational, then $\psi_{\ell}$ is $E^{\prime}$-rational.
Proof.
Apply Proposition 2.11 (by taking $N=[L:K]!$) and Proposition 2.12. ∎
2.8. $E$-rationality of abelian part
Given a semisimple $E$-rational $\ell$-adic representation
$\rho_{\ell}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell}),$
the abelian part of $\rho_{\ell}$, denoted by $\rho_{\ell}^{\text{ab}}$, is the maximal abelian subrepresentation of $\rho_{\ell}$. Suppose $\rho_{\ell}^{\text{ab}}$ is non-zero. Since $\rho_{\ell}^{\text{ab}}$ weakly divides $\rho_{\ell}$, it is locally algebraic by Theorem 1.2. Is it also $E$-rational? We conjecture the following and give some evidence (Theorem 2.15).
Conjecture 2.14.
If the algebraic monodromy group $\mathbf{G}_{\ell}$ of $\rho_{\ell}$ is connected and the abelian part $\rho_{\ell}^{\text{ab}}$ is non-zero, then $\rho_{\ell}^{\text{ab}}$ is $E$-rational.
Recall that the derived group of the identity component $\mathbf{G}_{\ell}^{\circ}$ is denoted by $\mathbf{G}_{\ell}^{\operatorname{der}}$; the group $\mathbf{G}_{\ell}^{\operatorname{der}}$ is semismiple. Decompose the representation $\rho_{\ell}=\rho_{\ell}^{\text{c}}\oplus\rho_{\ell}^{\text{ab}}$ and let $V_{\ell}^{\text{c}}$ (resp. $V_{\ell}^{\text{ab}}$) be the representation space of $\rho_{\ell}^{\text{c}}$ (resp. $\rho_{\ell}^{\text{ab}}$). Since $\mathbf{G}_{\ell}^{\operatorname{der}}$ acts trivially on $V_{\ell}^{\text{ab}}$, identify $\mathbf{G}_{\ell}^{\operatorname{der}}$ as a subgroup of $\mathrm{GL}_{V_{\ell}^{\text{c}}}$.
Theorem 2.15.
Let $\rho_{\ell}$ be a semisimple $E$-rational $\ell$-adic representation such that $V_{\ell}^{\text{c}}$ and $V_{\ell}^{\text{ab}}$ are both non-zero. If $\mathbf{G}_{\ell}$ is connected and the formal character of $\mathbf{G}_{\ell}^{\operatorname{der}}$ on $V_{\ell}^{\text{c}}$ does not have zero as a weight, then $\rho_{\ell}^{\text{ab}}$ is $E$-rational.
Proof.
Decompose $\rho_{\ell}^{\text{ab}}$ as a direct sum of characters:
(22) $\rho_{\ell}^{\text{ab}}=\chi_{1}\oplus\chi_{2}\oplus\cdots\oplus\chi_{m}.$
[…]
$\phi_{0}:\mathbf{S}_{\mathfrak{m}}\to\mathrm{GL}_{m,E^{\prime}}.$
Let $E_{\lambda}$ be the closure of $E$ in $\overline{\mathbb{Q}}_{\ell}$ (i.e., $\lambda\in\Sigma_{E}$). Then $E^{\prime}\otimes_{E}E_{\lambda}=\prod_{\lambda^{\prime}|\lambda}E^{\prime}_{\lambda^{\prime}}$ and by restriction of scalars [BGP19, Definition 3.4], we define
(23) $\phi_{\lambda}:\operatorname{Gal}_{K}^{\text{ab}}\stackrel{{\scriptstyle\epsilon_{\ell}}}{{\rightarrow}}\mathbf{S}_{\mathfrak{m}}(E_{\lambda})\stackrel{{\scriptstyle\prod_{\lambda^{\prime}|\lambda}\phi_{0}\times E^{\prime}_{\lambda^{\prime}}}}{{\longrightarrow}}\mathrm{GL}_{n}(E^{\prime}\otimes_{E}E_{\lambda})\subset\mathrm{GL}_{n[E^{\prime}:E]}(E_{\lambda})\subset\mathrm{GL}_{n[E^{\prime}:E]}(\overline{\mathbb{Q}}_{\ell}),$
which is $E$-rational and contains $\rho_{\ell}^{\text{ab}}$ as subrepresentation.
Let $L$ be a finite extension of $K$ such that the algebraic monodromy group $\mathbf{H}_{\ell}$ of the restriction $(\rho_{\ell}\oplus\phi_{\lambda})|_{\operatorname{Gal}_{L}}$ is connected. Then the Chebotarev density theorem implies the existence of a density one subset $\mathcal{L}_{L}\subset\Sigma_{L}$ such that for any distinct pair of characters $\chi,\chi^{\prime}$ of $\phi_{\lambda}|_{\operatorname{Gal}_{L}}$, one has
[…]
If $\rho_{\ell}^{\text{ab}}$ is not $E$-rational, then $\rho_{\ell}^{\text{ab}}|_{\operatorname{Gal}_{L}}$ is not $E$-rational by Corollary 2.13. Assume $\rho_{\ell}^{\text{ab}}|_{\operatorname{Gal}_{L}}$ does not contain any $E$-rational subrepresentation $\psi_{\ell}$, otherwise we can reduce to smaller $m$ (by passing to the quotient mod $\psi_{\ell}$) and use the induction hypothesis. Suppose
(25) $\chi_{1}|_{\operatorname{Gal}_{L}},\chi_{2}|_{\operatorname{Gal}_{L}},....,\chi_{k}|_{\operatorname{Gal}_{L}}$
is the set of distinct characters in $\rho_{\ell}^{\text{ab}}|_{\operatorname{Gal}_{L}}$. By property (24), the $k$ numbers
(26) $\chi_{1}|_{\operatorname{Gal}_{L}}(\operatorname{Frob}_{w}),\chi_{2}|_{\operatorname{Gal}_{L}}(\operatorname{Frob}_{w}),....,\chi_{k}|_{\operatorname{Gal}_{L}}(\operatorname{Frob}_{w})\in E^{\prime}$
are distinct for any $w\in\mathcal{L}_{L}$. Since the direct sum of the (locally algebraic) characters in (25) is not $E$-rational by assumption, Proposition 2.12 implies that for some $w\in\mathcal{L}_{L}$ the distinct numbers in (26) are not invariant under $\operatorname{Gal}(E^{\prime}/E)$. Hence, the $E$-rational $\phi_{\lambda}|_{\operatorname{Gal}_{L}}$ has some one-dimensional subrepresentation $\tau$ distinct from (25).
For almost all $w\in\Sigma_{L}$, the definition and the $E$-rationality of (23) imply that $\tau(\operatorname{Frob}_{w})$ must be a $\operatorname{Gal}(E^{\prime}/E)$-conjugate of some eigenvalue of $\rho_{\ell}^{\text{ab}}|_{\operatorname{Gal}_{L}}(\operatorname{Frob}_{w})$. As $\tau$ is distinct from (25), property (24) and $E$-rationality of $\rho_{\ell}|_{\operatorname{Gal}_{L}}=\rho_{\ell}^{\text{c}}|_{\operatorname{Gal}_{L}}\oplus\rho_{\ell}^{\text{ab}}|_{\operatorname{Gal}_{L}}$ imply that $\tau(\operatorname{Frob}_{w})$ is a root of the characteristic polynomial of $\rho_{\ell}^{\text{c}}|_{\operatorname{Gal}_{L}}(\operatorname{Frob}_{w})$ for all $w\in\mathcal{L}_{L}$.
Consider the twisted $\ell$-adic representation $\rho_{\ell}^{\text{c}}|_{\operatorname{Gal}_{L}}\otimes\tau^{-1}$ and let $\mathbf{G}_{\tau}$ be its algebraic monodromy group. The faithful representations of (the derived) $\mathbf{G}_{\ell}^{\operatorname{der}}$ and $\mathbf{G}_{\tau}^{\operatorname{der}}$ on respectively the representation spaces $V_{\ell}^{\text{c}}$ and $V_{\ell}^{\text{c}}\otimes\tau^{-1}$ are identified (as twisting by a character will not affect the derived group). Thus, the formal character of $\mathbf{G}_{\tau}^{\operatorname{der}}$ does not have zero weight (by assumption), which implies that the following conjugation-invariant subvariety of $\mathbf{G}_{\tau}$ is proper:
$\mathbf{Y}_{\tau}:=\{g\in\mathbf{G}_{\tau}:~{}1~{}\text{is an eigenvalue of}~{}g\}.$
This is impossible because $\rho_{\ell}^{\text{c}}|_{\operatorname{Gal}_{L}}\otimes\tau^{-1}(\operatorname{Frob}_{w})\in\mathbf{Y}_{\tau}$ for all $w\in\mathcal{L}_{L}$ and $\mathcal{L}_{L}\subset\Sigma_{L}$ is a density one subset. ∎
We prove Conjecture 2.14 when $n=3$.
Corollary 2.16.
Let $\rho_{\ell}$ be a three-dimensional semisimple $E$-rational $\ell$-adic representation with non-zero $\rho_{\ell}^{\text{ab}}$. If $\mathbf{G}_{\ell}$ is connected, then $\rho_{\ell}^{\text{ab}}$ is $E$-rational.
Proof.
We only need to handle the case that $\rho_{\ell}^{\text{ab}}$ is one-dimensional. Then $\mathbf{G}_{\ell}^{\operatorname{der}}$ is $\mathrm{SL}_{2}$ on $V_{\ell}^{\text{c}}$, which has no zero weight in the formal character. We are done by Theorem 2.15. ∎
2.9. Abelian and weak abelian part of compatible system
Let $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda})\}$ be a semisimple Serre compatible system of $K$ defined over $E$ unramified outside a finite $S\subset\Sigma_{K}$. Let $\mathbf{G}_{\lambda}$ be the algebraic monodromy group of $\rho_{\lambda}$. The connectedness and formal character of $\mathbf{G}_{\lambda}$ is independent of $\lambda$; see [Se81b]. We say that $\{\rho_{\lambda}\}$ is connected if $\mathbf{G}_{\lambda}$ is connected for some $\lambda$ (hence all $\lambda$). Define $\{\rho_{\lambda}^{\text{ab}}\}$ (resp. $\{\rho_{\lambda}^{\text{wab}}\}$, well-defined if $\{\rho_{\lambda}\}$ is connected by Proposition 2.9) as the abelian part (resp. weak abelian part) of the Serre compatible system $\{\rho_{\lambda}\}$. One can ask if $\{\rho_{\ell}^{\text{ab}}\}$ (resp. $\{\rho_{\lambda}^{\text{wab}}\}$) is still a Serre compatible system defined over $E$ when it is non-zero. This question is studied in [Hu20] (in case $E=\mathbb{Q}$) for the abelian part by using some $\lambda$-independence results of $\mathbf{G}_{\lambda}$, such as the following.
Theorem 2.17.
[Hu13, Theorem 3.19, Remark 3.22] The formal bi-character of $\mathbf{G}_{\lambda}$ is independent of $\lambda$. In particular, the formal character of $\mathbf{G}_{\lambda}$ (resp. $\mathbf{G}_{\lambda}^{\operatorname{der}}$) is independent of $\lambda$.
By Theorem 2.17, if $\mathbf{G}_{\lambda_{0}}^{\operatorname{der}}=\mathrm{SL}_{n}$ (resp. $\mathbf{G}_{\lambda_{0}}=\mathrm{GL}_{n}$) for some $\lambda_{0}$ then $\mathbf{G}_{\lambda}^{\operatorname{der}}=\mathrm{SL}_{n}$ (resp. $\mathbf{G}_{\lambda}=\mathrm{GL}_{n}$) for all $\lambda$. Thus, we obtain $\lambda$-independence of $\mathbf{G}_{\lambda}^{\operatorname{der}}$ (resp. $\mathbf{G}_{\lambda}$). However when $n=3$ and $\mathbf{G}_{\lambda_{0}}^{\operatorname{der}}=\mathrm{SL}_{2}\times\{1\}\subset\mathrm{GL}_{3}$ for some $\lambda_{0}$, the semisimple $\mathbf{G}_{\lambda}^{\operatorname{der}}$ is either $\mathrm{SL}_{2}\times\{1\}$ or $\mathrm{SO}_{3}$ by Theorem 2.17 and we need extra information to exclude the second possibility.
Proposition 2.18.
Let $\{\rho_{\lambda}\}$ be a connected semisimple three-dimensional Serre compatible system of $K$ defined over $E$. Suppose $\rho_{\lambda_{0}}=\sigma_{\lambda_{0}}\oplus\tau_{\lambda_{0}}$ is a decomposition into irreducible factors for some $\lambda_{0}$ such that $\sigma_{\lambda_{0}}$ is two-dimensional. Then the following assertions hold.
  1. (i)
    $\tau_{\lambda_{0}}$ satisfies ($E$-SCS), i.e., $\tau_{\lambda_{0}}$ belongs to a SCS $\{\tau_{\lambda}\}$ defined over $E$.
  2. (ii)
    If $\rho_{\lambda}$ is reducible, then $\mathbf{G}_{\lambda}^{\operatorname{der}}=\mathrm{SL}_{2}\times\{1\}$ and the (unique) one-dimensional subrepresentation is $\tau_{\lambda}$.
  3. (iii)
    If $\rho_{\lambda}$ is irreducible, then $\mathbf{G}_{\lambda}^{\operatorname{der}}=\mathrm{SO}_{3}$ and $\det(\rho_{\lambda})=\tau_{\lambda}^{3}$.
Proof.
[…]
$\mathrm{SL}_{2}\times\{1\}\subset\mathbf{G}_{\lambda_{0}}^{\prime}\subset\mathrm{GL}_{2}\times\{1\},$
the formal character of $\mathbf{G}_{\lambda}^{\prime}$ has a zero weight for all $\lambda$ by Theorem 2.17. Hence, either $\mathrm{SL}_{2}\times\{1\}\subset\mathbf{G}_{\lambda}^{\prime}\subset\mathrm{GL}_{2}\times\{1\}$ or $\mathbf{G}_{\lambda}^{\prime}=\mathrm{SO}_{3}$, depending on the reducibility of $\rho_{\lambda}$. This implies the last parts in assertions (ii) and (iii). ∎
Corollary 2.19.
Let $\{\rho_{\lambda}\}$ be a connected semisimple three-dimensional Serre compatible system of $K$ defined over $E$. If $\rho_{\lambda}$ is reducible for all $\lambda$, then $\{\rho_{\lambda}^{\text{ab}}\}$ is a Serre compatible system defined over $E$.
Proof.
Let $\lambda_{0}$ be a finite place of $E$. If $\rho_{\lambda_{0}}$ is abelian, then the whole system is abelian by §$\mathsection 2.4$ and thus $\{\rho_{\lambda}^{\text{ab}}\}=\{\rho_{\lambda}\}$ is a compatible system. If $\rho_{\lambda_{0}}$ is non-abelian, then it is the direct sum of a two-dimensional irreducible representation and a character. Since $\rho_{\lambda}$ is reducible for all $\lambda$, the assertion follows from Proposition 2.18(i),(ii). ∎
Although the compatibility of the abelian part $\{\rho_{\lambda}^{\text{ab}}\}$ is not known, this is not the case for the weak abelian part $\{\rho_{\lambda}^{\text{wab}}\}$ by the following theorem. Since the formal character of $\mathbf{G}_{\lambda}^{\operatorname{der}}$ is independent of $\lambda$ (Theorem 2.17), the multiplicity of the weight zero in the formal character is independently of $\lambda$; we denote it by $n_{0}$. [^2] The independence of $n_{0}$ of $\lambda$ can also be deduced from the proof of Theorem 2.20 by choosing the first $\lambda$ in such a way that its $n_{0}$ is maximal.
Theorem 2.20.
Suppose $\{\rho_{\lambda}\}$ is connected and $n_{0}>0$. There exists a finite extension $E^{\prime}$ of $E$ such that if $\{\rho_{\lambda}\}$ is considered as a Serre compatible system $\{\rho_{\lambda^{\prime}}\}$ of $K$ defined over $E^{\prime}$ [^3] We define $\rho_{\lambda^{\prime}}:=\rho_{\lambda}$ if $\lambda^{\prime}\in\Sigma_{E^{\prime}}$ divides $\lambda\in\Sigma_{E}$., then the weak abelian part $\{\rho_{\lambda^{\prime}}^{\text{wab}}\}$ is an $n_{0}$-dimensional Serre compatible system defined over $E^{\prime}$.
Proof.
Let $\lambda_{0}$ be a finite place of $E$. Since the weak abelian part $\rho_{\lambda_{0}}^{\text{wab}}$ of $\rho_{\lambda_{0}}$ is locally algebraic (Theorem 1.2), it can be extended to an abelian semisimple Serre compatible system $\{\psi_{\lambda^{\prime}}\}$ defined over some finite extension $E^{\prime}$ of $E$ such that $\rho_{\lambda_{0}}^{\text{wab}}=\psi_{\lambda_{0}^{\prime}}$. For $\lambda^{\prime}\in\Sigma_{E^{\prime}}$, since $\psi_{\lambda^{\prime}}$ is a degree $n_{0}$ weak abelian direct summand of $\rho_{\lambda^{\prime}}$, it coincides with $\rho_{\lambda^{\prime}}^{\text{wab}}$ by Proposition 2.9. We are done. ∎
3. Irreducibility and $\ell$-adic Hodge properties
Suppose $K$ is a totally real field (or CM) field. Let $\pi$ be a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{n}(\mathbb{A}_{K})$. For a rational prime $\ell$, denote by $S_{\ell}$ the set of finite places $v\in\Sigma_{K}$ above $\ell$.
3.1. Serre compatible system attached to $\pi$
By Clozel [Cl90, Theorem 3.13, Lemmas 3.14, 3.15], $\pi$ is C-arithmetic. By [BG14, Proposition 5.2.3], $\pi$ is C-arithmetic if and only if $\pi\otimes|\det|^{\frac{1-n}{2}}$ is L-arithmetic. Thus, there exist a number field $E$ and a finite subset $S\subset\Sigma_{K}$ containing the ramified primes of $\pi$ such that the Satake parameters of $\pi_{v}\otimes|\det|_{v}^{\frac{1-n}{2}}$ are defined over $E$ for all $v\in\Sigma_{K}\backslash S$. Any field isomorphism $\iota:\mathbb{C}\to\overline{\mathbb{Q}}_{\ell}$ induces a finite place $\lambda\in\Sigma_{E}$ dividing $\ell$, so that $\overline{\mathbb{Q}}_{\ell}$ can be identified as $\overline{E}_{\lambda}$. Then the semisimple $\ell$-adic representation (1) of $K$,
(27) $\rho_{\pi,\iota}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{\mathbb{Q}}_{\ell})=\mathrm{GL}_{n}(\overline{E}_{\lambda}),$
is $E$-rational and unramified outside $S$. If $\iota^{\prime}:\mathbb{C}\to\overline{\mathbb{Q}}_{\ell}$ is another isomorphism that induces the same $\lambda\in\Sigma_{E}$, then the semisimple $\rho_{\pi,\iota}$ and $\rho_{\pi,\iota^{\prime}}$ are equivalent representation by (2), which we denote by $\rho_{\lambda}$. Hence, the set $\{\rho_{\pi,\iota}:~{}\iota:\mathbb{C}\to\overline{\mathbb{Q}}_{\ell},~{}\ell~{}\text{rational prime}\}$ of Galois representations is equal to the family of $\ell$-adic representations,
(28) $\{\rho_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{n}(\overline{E}_{\lambda}):~{}\lambda\in\Sigma_{E}\},$
which by (2), is a semisimple Serre compatible system of $K$ defined over $E$ unramified outside $S$. Let $\mathbf{G}_{\lambda}$ be the algebraic monodromy group of $\rho_{\lambda}$.
3.2. Irreducibility and algebraic monodromy groups
Suppose $K$ is totally real and $n=3$. We prove the irreducibility of $\rho_{\lambda}$ (Theorem 1.1) by an L-function argument in [Ram13] and study the $\lambda$-independence of $\mathbf{G}_{\lambda}$.
3.2.1. Proof of Theorem 1.1.
If $\rho_{\lambda}$ is reducible, we have a decomposition $\rho_{\lambda}=\sigma_{\lambda}\oplus\tau_{\lambda}$ where $\sigma_{\lambda}$ is two-dimensional and $\tau_{\lambda}$ is a character. Then
(29) $\begin{split}\mathrm{Alt}^{2}(\rho_{\lambda})&=(\sigma_{\lambda}\otimes\tau_{\lambda})\oplus\det(\sigma_{\lambda}),\\ \tau_{\lambda}^{2}\oplus\mathrm{Alt}^{2}(\rho_{\lambda})&=(\rho_{\lambda}\otimes\tau_{\lambda})\oplus\det(\sigma_{\lambda}),\\ 1\oplus(\mathrm{Alt}^{2}(\rho_{\lambda})\otimes\tau_{\lambda}^{-2})&=(\rho_{\lambda}\otimes\tau_{\lambda}^{-1})\oplus(\det(\sigma_{\lambda})\cdot\tau_{\lambda}^{-2}).\end{split}$
Since $\tau_{\lambda}$ is locally algebraic by Theorem 1.2, it comes from an algebraic Hecke character $\tau$ of $K$. Since $\rho_{\lambda}$ is unramified outside $S^{\prime}=S\cup S_{\ell}$, we obtain from (29) and the local-global compatibility (2) an equation on partial L-functions
(30) $L^{S^{\prime}}(1,s)L^{S^{\prime}}(\mathrm{Alt}^{2}(\pi)\otimes\psi_{1},s)=L^{S^{\prime}}(\pi\otimes\psi_{2},s)L^{S^{\prime}}(\psi_{3},s),$
where $\psi_{1},\psi_{2},\psi_{3}$ are some algebraic Hecke characters of $K$ and $\psi_{3}$ corresponds to the locally algebraic character $\det(\sigma_{\lambda})\cdot\tau_{\lambda}^{-2}$.
Suppose $\det(\sigma_{\lambda})\cdot\tau_{\lambda}^{-2}$ is non-trivial. On the one hand, the right-hand side of (30) at $s=1$ belongs to $\mathbb{C}$ by $\psi_{3}$ non-trivial and $\pi$ cuspidal. On the other hand, $L^{S^{\prime}}(1,s)$ has a simple pole at $s=1$ and $L^{S^{\prime}}(\mathrm{Alt}^{2}(\pi)\otimes\psi_{1},s)$ is non-zero at $s=1$ by Shahidi [Sh97, Theorem 1.1]. We get a contradiction.
Suppose $\det(\sigma_{\lambda})\cdot\tau_{\lambda}^{-2}$ is trivial. Then $\rho_{\lambda}\otimes\tau_{\lambda}^{-1}$ is self-dual. Since $\tau_{\lambda}$ is locally algebraic, $\pi$ is essentially self-dual by multiplicity one [JS81]. But in this case, $\rho_{\lambda}$ is irreducible (see [Hu23a, §$\mathsection 4.4.3$]). We conclude that $\rho_{\lambda}$ is irreducible.
3.2.2. Algebraic monodromy groups
The faithful representation $\mathbf{G}_{\lambda}\hookrightarrow\mathrm{GL}_{3,\overline{E}_{\lambda}}$ is irreducible for all $\lambda$ by Theorem 1.1. If $\rho_{\lambda}$ is Lie-irreducible [^4] The faithful representation $\mathbf{G}_{\lambda}^{\circ}\to\mathrm{GL}_{3,\overline{E}_{\lambda}}$ is irreducible., there are two types:
  1. ($A_{1}$):
    $\mathrm{SO}_{3,\overline{E}_{\lambda}}\subset\mathbf{G}_{\lambda}\subset\mathrm{GO}_{3,\overline{E}_{\lambda}}$;
  2. ($A_{2}$):
    $\mathrm{SL}_{3,\overline{E}_{\lambda}}\subset\mathbf{G}_{\lambda}$.
If $\rho_{\lambda}$ is not Lie-irreducible, there are two types by [Cl37, Theorems 1,2] (see also [CG13, Lemma 4.3]):
  1. (D):
    $\rho_{\lambda}(\operatorname{Gal}_{L})$ belongs to the center of $\mathrm{GL}_{3,\overline{E}_{\lambda}}$ for some finite extension $L/K$;
  2. (Ind):
    $\rho_{\lambda}=\mathrm{Ind}^{K}_{L}\psi_{\lambda}$ is induced from a locally algebraic character $\psi_{\lambda}:\operatorname{Gal}_{L}\to\overline{E}_{\lambda}^{\times}$ for some cubic extension $L/K$ and the three characters of the abelian $\rho_{\lambda}|_{\operatorname{Gal}_{L^{\prime}}}$ are distinct for some $L^{\prime}/L$.
Note that $\mathbf{G}_{\lambda}^{\operatorname{der}}$ is trivial for type (D) and (Ind).
Theorem 3.1.
The irreducible type of $\rho_{\lambda}$ is independent of $\lambda$. Moreover, the following assertion hold.
  1. (i)
    Suppose $\rho_{\lambda}$ is of type ($A_{1}$). The automorphic representation $\pi$ is essentially self-dual.
  2. (ii)
    Suppose $\rho_{\lambda}$ is of type (Ind). After replacing $E$ by a finite extension, the compatible system $\{\rho_{\lambda}\}_{\lambda}$ of $K$ is induced from a one-dimensional $E$-rational compatible system $\{\psi_{\lambda}\}_{\lambda}$ of a cubic extension $L/K$.
Proof.
Let $\lambda_{0}\in\Sigma_{E}$. If $\rho_{\lambda_{0}}$ is of type ($A_{1}$) (resp. ($A_{2}$)), then $\mathbf{G}_{\lambda_{0}}^{\operatorname{der}}=\mathrm{SO}_{3}$ (resp. $\mathrm{SL}_{3}$). Irreducibility of $\rho_{\lambda}$ and $\lambda$-independence of the formal character of $\mathbf{G}_{\lambda}^{\operatorname{der}}$ (Theorem 2.17) imply that $\mathbf{G}_{\lambda}^{\operatorname{der}}=\mathrm{SO}_{3}$ (resp. $\mathrm{SL}_{3}$). Moreover, the $E$-rational adjoint representation $\operatorname{ad}(\rho_{\lambda_{0}}):=\rho_{\lambda_{0}}\otimes\rho_{\lambda_{0}}^{\vee}$ is the sum of an irreducible eight-dimensional representation and a locally algebraic character $\tau_{\lambda_{0}}=\operatorname{ad}(\rho_{\lambda_{0}})^{\text{ab}}$ (Theorem 1.2) such that $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\tau_{\lambda_{0}}.$ Hence, $\pi$ is essentially self-dual by multiplicity one [JS81].
If $\rho_{\lambda_{0}}$ is of type (D), the adjoint representation $\operatorname{ad}(\rho_{\lambda_{0}})=\rho_{\lambda_{0}}\otimes\rho_{\lambda_{0}}^{\vee}$ has finite image. Since $\{\operatorname{ad}(\rho_{\lambda}):=\rho_{\lambda}\otimes\rho_{\lambda}^{\vee}\}$ is a Serre compatible system, $\operatorname{ad}(\rho_{\lambda})$ has finite image for all $\lambda$ by [Se81b], which implies $\rho_{\lambda}(\operatorname{Gal}_{L})$ belong to the homothety for all $\lambda$.
If $\rho_{\lambda_{0}}$ is of type (Ind), then it is potentially abelian and $E$-rational, which imply that $\psi_{\lambda_{0}}$ is $E^{\prime}$-rational for some $E\subset E^{\prime}\subset\overline{E}_{\lambda}$. There exists a Serre compatible system $\{\psi_{\lambda^{\prime}}:~{}\lambda^{\prime}\in\Sigma_{E^{\prime}}\}$ of $L$ defined over $E^{\prime}$ extending $\psi_{\lambda_{0}}$ and moreover, $\{\mathrm{Ind}^{L}_{K}\psi_{\lambda^{\prime}}\}$ is a semisimple Serre compatible system of $K$ containing $\{\rho_{\lambda}\}$. Since $\rho_{\lambda_{0}}|_{\operatorname{Gal}_{L^{\prime}}}$ is a direct sum of three distinct locally algebraic characters, this is true for $\rho_{\lambda}|_{\operatorname{Gal}_{L^{\prime}}}$ for all $\lambda$ by compatibility. We are done. ∎
Remark 3.2.
The arguments in this section do not use the regularity of $\pi$. In next subsection, this condition will be used to prove that $\rho_{\lambda}$ cannot be of type (D), see Proposition 3.5.
3.3. $\ell$-adic Hodge properties
Suppose $K$ is totally real and $\pi$ is a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{3}(\mathbb{A}_{K})$. The weight $(a_{\sigma,i})\in(\mathbb{Z}^{3})^{\text{Hom}(K,\mathbb{C})}$ of $\pi$ (see [BLGGT14b, §$\mathsection 2.1$]) satisfies
$a_{\sigma,1}\geq a_{\sigma,2}\geq a_{\sigma,3}$
for all $\sigma\in\text{Hom}(K,\mathbb{C})$. Let $\{\rho_{\lambda}:=\rho_{\pi,\iota_{\lambda}}\}$ be the SCS defined over $E$ attached to $\pi$ in (28), where $\iota_{\lambda}:\mathbb{C}\to\overline{E}_{\lambda}$ is an isomorphism for all $\lambda$. Recall that the residue characteristic of $\lambda\in\Sigma_{E}$ is denoted $\ell(\lambda)$.
By Theorem 3.1, we define the irreducible type of the SCS $\{\rho_{\lambda}\}$ as the type of $\rho_{\lambda}$. If $\{\rho_{\lambda}\}$ is of type (D) or (Ind), then the SCS is potentially abelian and thus $\rho_{\lambda}$ is de Rham for all $\lambda$. If $\{\rho_{\lambda}\}$ is of type ($A_{1}$), then $\pi$ is essentially self-dual (Theorem 3.1(i)) and by [BLGGT12, BLGGT14a] and [Ca14], the following $\ell$-adic Hodge properties hold for $v\in S_{\ell(\lambda)}$.
  1. (dR):
    The local representation $\rho_{\lambda}|_{\operatorname{Gal}_{K_{v}}}$ is de Rham with (distinct) $\tau$-Hodge-Tate numbers
    $HT_{\tau}:=\{a_{\iota_{\lambda}^{-1}\circ\tau}+2,~{}a_{\iota_{\lambda}^{-1}\circ\tau}+1,~{}a_{\iota_{\lambda}^{-1}\circ\tau}\},$
    where $\tau:K_{v}\to\overline{E}_{\lambda}$.
  2. (Cr):
    If $\pi_{v}$ is unramified, then $\rho_{\lambda}|_{\operatorname{Gal}_{K_{v}}}$ is crystalline.
Denote by $\mathbb{F}_{\lambda}$ the residue field of $E_{\lambda}$ and by $\bar{\rho}_{\lambda}:\operatorname{Gal}_{K}\to\mathrm{GL}_{3}(\overline{\mathbb{F}}_{\lambda})$ the residual representation given by the semisimplified reduction of $\rho_{\lambda}$. A recent result of A’Campo [A’C23, Theorem 1.0.6] asserts that the $\ell$-adic Hodge properties (dR) and (Cr) hold if
  1. (RI):
    $\rho_{\lambda}$ is residually irreducible (i.e., $\bar{\rho}_{\lambda}$ is irreducible) and
  2. (DG):
    $\bar{\rho}_{\lambda}$ is decomposed generic (see [ACC+23, Definition 4.3.1]).
A sufficient condition of (DG) is as follows.
Lemma 3.3.
[ACC+23, Lemma 7.1.5] Suppose $L$ is a number field and $\bar{r}:\operatorname{Gal}_{L}\to\mathrm{GL}_{n}(\overline{\mathbb{F}}_{\ell})$ is a continuous representation. Denote by $\operatorname{ad}(\bar{r})$ the adjoint representation $\bar{r}\otimes\bar{r}^{\vee}$ and by $M$ the normal closure of $\overline{L}^{\mathrm{Ker}(\operatorname{ad}(\bar{r}))}$ over $\mathbb{Q}$. If $M$ does not contain a primitive $\ell$th root of unity, then $\bar{r}$ is decomposed generic.
Remark 3.4.
Conversely, if $\{\rho_{\lambda}\}$ is a strictly compatible system in the sense of [BLGGT14b, Section 5.1] with (distinct) $\tau$-Hodge-Tate weights (e.g., when $\{\rho_{\lambda}\}$ satisfies the local-global compatibility (2) for all $v\in\Sigma_{K}$ and also (dR) and (Cr) above), then (RI) holds for almost all $\lambda$ by Theorem 1.1 and [Hu23b, Theorem 1.1(iii)].
We are now ready to exclude irreducible type (D).
Proposition 3.5.
The compatible system $\{\rho_{\lambda}\}$ cannot be of type (D).
Proof.
Suppose $\{\rho_{\lambda}\}$ is of type (D). Consider the SCS $\{\operatorname{ad}(\rho_{\lambda}):=\rho_{\lambda}\otimes\rho_{\lambda}^{\vee}\}$. By Serre [Se81b], the finite $\mathrm{Ker}(\operatorname{ad}(\rho_{\lambda}))$ is independent of $\lambda$ and is equal to $\operatorname{Gal}(K^{\prime}/K)$. Let $M$ be the normal closure of $K^{\prime}$ over $\mathbb{Q}$ and $m:=[M:\mathbb{Q}]$.
Pick $\lambda\in\Sigma_{E}$ such that $\ell:=\ell(\lambda)>1+m$. Since the normal closure of $\overline{K}^{\mathrm{Ker}(\operatorname{ad}(\bar{\rho}_{\lambda}))}$ over $\mathbb{Q}$ is also equal to $M$ and $M$ cannot contain a primitive $\ell$th root of unity, Lemma 3.3 implies that $\bar{\rho}_{\lambda}$ is decomposed generic (DG). Moreover as $\ell>m\geq|\operatorname{ad}(\bar{\rho}_{\lambda})|$, [Se77, Proposition 43(ii)] asserts that the irreducible factors of the semisimple representations $\operatorname{ad}(\rho_{\lambda})$ and $\operatorname{ad}(\bar{\rho}_{\lambda})$ correspond via reduction. Since the multiplicity of the trivial representation of $\operatorname{ad}(\rho_{\lambda})$ is one (by the irreducibility of $\rho_{\lambda}$), this also holds for $\operatorname{ad}(\bar{\rho}_{\lambda})$, which implies that $\bar{\rho}_{\lambda}$ is irreducible (RI). Therefore, [A’C23, Theorem 1.0.6] asserts that $\rho_{\lambda}|_{\operatorname{Gal}_{K_{v}}}$ (for $v\in S_{\ell}$) is de Rham with distinct $\tau$-Hodge-Tate numbers. But this contradicts that $\rho_{\lambda}(\operatorname{Gal}_{L})$ belongs to the homothety for some finite extension $L/K$. ∎
It remains to study the $\ell$-adic Hodge properties of $\{\rho_{\lambda}\}$ when it is of type ($A_{2}$).
Theorem 3.6.
Suppose $\{\rho_{\lambda}\}$ is of type ($A_{2}$). Then there exists a Dirichlet density one set $\mathcal{L}\subset\Sigma_{\mathbb{Q}}$ of primes $\ell$ such that $\rho_{\lambda}$ satisfies the above properties (dR) and (Cr) if $\ell\in\mathcal{L}$.
Proof.
By [A’C23, Theorem 1.0.6] and Lemma 3.3, it suffices to show that for a Dirichlet density one set of primes $\ell$, the following two assertions hold whenever $\ell(\lambda)=\ell$:
  1. (RI):
    the residual representation $\bar{\rho}_{\lambda}$ is irreducible and
(31) $\mathbb{Q}(\zeta_{\ell})\nsubseteq M_{\lambda}$
where $\zeta_{\ell}$ is a primitive $\ell$th root of unity and $M_{\lambda}$ is the normal closure of $\overline{K}^{\mathrm{Ker}(\operatorname{ad}(\bar{\rho}_{\lambda}))}$ over $\mathbb{Q}$.
[…]
  • $\{\mathrm{Tr}(\rho_{\lambda}(\gamma)):~{}\gamma\in\operatorname{Gal}_{K}\}\subset E_{\lambda}$ and
  • there exists $\gamma\in\operatorname{Gal}_{K}$ such that $\rho_{\lambda}(\gamma)$ has three distinct roots in $E$.
Hence, we may assume $\rho_{\lambda}(\operatorname{Gal}_{K})\subset\mathrm{GL}_{3}(E_{\lambda})$ for all $\lambda$ by [BLGGT14b, Lemma A.1.5].
Consider the following semisimple Serre compatible systems (defined over $\mathbb{Q}$ unramified outside $S$) given by restriction of scalars:
(32) $\{\Phi_{\ell}:\operatorname{Gal}_{K}\stackrel{{\scriptstyle\prod_{\lambda|\ell}\rho_{\lambda}}}{{\longrightarrow}}\prod_{\lambda|\ell}\mathrm{GL}_{3}(E_{\lambda})=\mathrm{GL}_{3}(E\otimes\mathbb{Q}_{\ell})=\mathrm{Res}_{E/\mathbb{Q}}(\mathrm{GL}_{3,E})(\mathbb{Q}_{\ell})\subset\mathrm{GL}_{3[E:\mathbb{Q}]}(\mathbb{Q}_{\ell})\}_{\ell}$
and
(33) $\{\Phi_{\ell}^{\operatorname{ad}}:\operatorname{Gal}_{K}\stackrel{{\scriptstyle\prod_{\lambda|\ell}\operatorname{ad}(\rho_{\lambda})}}{{\longrightarrow}}\prod_{\lambda|\ell}\mathrm{GL}_{9}(E_{\lambda})=\mathrm{GL}_{9}(E\otimes\mathbb{Q}_{\ell})=\mathrm{Res}_{E/\mathbb{Q}}(\mathrm{GL}_{9,E})(\mathbb{Q}_{\ell})\subset\mathrm{GL}_{9[E:\mathbb{Q}]}(\mathbb{Q}_{\ell})\}_{\ell}.$
Let $\mathbf{H}_{\ell}$ be the algebraic monodromy group of $\Phi_{\ell}$. By Serre [Se81b], one can pick a finite extension $L$ of $K$ that is Galois over $\mathbb{Q}$ such that $\Phi_{\ell}(\operatorname{Gal}_{L})\subset\mathbf{H}_{\ell}^{\circ}$ for all $\ell$. It is enough to show that for a density one set $\mathcal{L}$ of rational primes $\ell$, the following two assertions hold whenever $\ell(\lambda)=\ell$:
  1. (RIL):
    the residual representation $\bar{\rho}_{\lambda}|_{\operatorname{Gal}_{L}}$ is (absolutely) irreducible and
(34) $\mathbb{Q}(\zeta_{\ell})\nsubseteq\Omega_{\ell},$
where $\Omega_{\ell}$ is the normal closure of $\overline{L}^{\mathrm{Ker}(\Phi_{\ell}^{\operatorname{ad}}|_{\operatorname{Gal}_{L}})}$ over $\mathbb{Q}$. Note that $\Omega_{\ell}/\mathbb{Q}$ is an infinite extension.
3.3.1. Proof of (RIL) for density one
Let $\mathbf{H}_{\ell}^{\operatorname{ad}}$ be the algebraic monodromy group of $\Phi_{\ell}^{\operatorname{ad}}|_{\operatorname{Gal}_{L}}$. Since $\Phi_{\ell}^{\operatorname{ad}}|_{\operatorname{Gal}_{L}}$ factors through $\Phi_{\ell}|_{\operatorname{Gal}_{L}}$ for all $\ell$ and the algebraic monodromy group of $\operatorname{ad}(\rho_{\lambda})$ is $\mathrm{PGL}_{3}$ for all $\lambda$, $\mathbf{H}_{\ell}^{\operatorname{ad}}$ is isomorphic to the adjoint quotient of $\mathbf{H}_{\ell}^{\circ}$, which is connected and adjoint semisimple.
Let $\alpha_{\ell}:\mathbf{H}_{\ell}^{\operatorname{sc}}\to\mathbf{H}_{\ell}^{\operatorname{ad}}$ be the universal covering. By Larsen [La95, Theorem 3.17], there is a Dirichlet density one set $\mathcal{L}^{\prime}$ of primes $\ell$ such that the preimage $\Gamma_{\ell}^{\operatorname{sc}}$ of $\Gamma_{\ell}^{\operatorname{ad}}:=\Phi_{\ell}^{\operatorname{ad}}(\operatorname{Gal}_{L})$ in $\mathbf{H}_{\ell}^{\operatorname{sc}}(\mathbb{Q}_{\ell})$ is a hyperspecial maximal compact subgroup for all $\ell\in\mathcal{L}^{\prime}$. It follows from [HL20, §$\mathsection 4.2$ Proof of Theorem 1.2] that if $\ell\in\mathcal{L}^{\prime}$ is sufficiently large, the following equality on invariant dimensions hold:
(35) $\dim_{\mathbb{Q}_{\ell}}\operatorname{ad}(\Phi_{\ell}^{\operatorname{ad}})^{\operatorname{Gal}_{L}}=\dim_{\mathbb{F}_{\ell}}\operatorname{ad}(\bar{\Phi}_{\ell}^{\operatorname{ad}})^{\operatorname{Gal}_{L}},$
but this implies
(36) $\dim_{\mathbb{Q}_{\ell}}(\Phi_{\ell}^{\operatorname{ad}})^{\operatorname{Gal}_{L}}=\dim_{\mathbb{F}_{\ell}}(\bar{\Phi}_{\ell}^{\operatorname{ad}})^{\operatorname{Gal}_{L}}$
since $\Phi_{\ell}^{\operatorname{ad}}$ is a subrepresentation of $\operatorname{ad}(\Phi_{\ell}^{\operatorname{ad}})$. Here $\bar{\Phi}_{\ell}^{\operatorname{ad}}$ is the (semisimple) residual representation of $\Phi_{\ell}^{\operatorname{ad}}$ and $\operatorname{ad}(\bar{\Phi}_{\ell}^{\operatorname{ad}})$ (resp. $\operatorname{ad}(\bar{\Phi}_{\ell}^{\operatorname{ad}})|_{\operatorname{Gal}_{L}}$ as $L/K$ Galois) is semisimple [Se94].
As $\{\rho_{\lambda}\}$ is of type ($A_{2}$), $\rho_{\lambda}|_{\operatorname{Gal}_{L}}$ is absolutely irreducible for all $\lambda$. By (32) and (36), $\bar{\rho}_{\lambda}|_{\operatorname{Gal}_{L}}$ is absolutely irreducible if $\ell$ belongs to a Dirichlet density one set $\mathcal{L}^{\prime\prime}$.
3.3.2. Proof of (34) for density one
We need some preparations before the proof.
Definition 3.7.
Let $G$ be a finite group and denote by $\mathrm{JH}(G)$ the multiset of Jordan-Hölder factors of $G$. Define $c_{\ell}(G)\in\mathbb{N}$ to be the product of the orders of the cyclic elements of $\mathrm{JH}(G)$ not isomorphic to $\mathbb{Z}/\ell\mathbb{Z}$.
Lemma 3.8.
Let $H_{1},H_{2},...,H_{k}$ be finite groups and $G$ a subgroup of $H_{1}\times H_{2}\times\cdots\times H_{k}$. If $G$ surjects onto $H_{i}$ for all $1\leq i\leq k$, then $\mathrm{JH}(G)\subset\mathrm{JH}(H_{1}\times H_{2}\times\cdots\times H_{k})$ as inclusion of multisets.
Proof.
For $1\leq i\leq k$, let $N_{i}$ be the kernel of $G$ projecting to $H_{1}\times H_{2}\times\cdots\times H_{i}$. Then we have a chain of normal subgroups (of $G$)
$\{e\}=N_{k}\subset\cdots\subset N_{2}\subset N_{1}\subset G$
such that $N_{i}/N_{i+1}$ injects into $H_{i+1}$. Since $G$ surjects onto $H_{i+1}$ and $N_{i}$ is normal in $G$, $N_{i}/N_{i+1}$ is normal in $H_{i+1}$. We are done. ∎
[…]
$\psi_{\ell}:\operatorname{Gal}(\Omega_{\ell}/L)\to\mathbb{Z}/(\ell-1)\mathbb{Z}.$
Let $L_{1}:=\overline{L}^{\text{Ker}(\Phi_{\ell}^{\operatorname{ad}}|_{\operatorname{Gal}_{L}})}$ and denote by
(37) $L_{1},L_{2},...,L_{k}$
be the $\operatorname{Gal}_{\mathbb{Q}}$-conjugates of $L_{1}$. Note that
  1. (i)
    $L_{1},L_{2}...,L_{k}$ generate $\Omega_{\ell}$;
  2. (ii)
    $k\leq[L:\mathbb{Q}]$;
  3. (iii)
    $\operatorname{Gal}(L_{i}/L)\cong\Gamma_{\ell}^{\operatorname{ad}}:=\Phi_{\ell}^{\operatorname{ad}}(\operatorname{Gal}_{L})$ for all $1\leq i\leq k$.
By looking at the action of $G_{\ell}:=\operatorname{Gal}(\Omega_{\ell}/L)$ on (37), we can identify it as a subgroup of
$\prod_{i=1}^{k}\operatorname{Gal}(L_{i}/L)\simeq\prod_{i=1}^{k}\Gamma_{\ell}^{\operatorname{ad}}$
that surjects onto the $i$th factor $\Gamma_{\ell}^{\operatorname{ad}}$ for all $i$. Identify $\Gamma_{\ell}^{\operatorname{ad}}$ as a subgroup of $\mathrm{GL}_{9[E:\mathbb{Q}]}(\mathbb{Z}_{\ell})$ using (33) and let $\bar{\Gamma}_{\ell}^{\operatorname{ad}}$ be the residual image in $\mathrm{GL}_{9[E:\mathbb{Q}]}(\mathbb{F}_{\ell})$. Let $\bar{G}_{\ell}$ be the image of $G_{\ell}$ in
(38) $\prod_{i=1}^{k}\bar{\Gamma}_{\ell}^{\operatorname{ad}}.$
Since the kernel of $\mathrm{GL}_{9[E:\mathbb{Q}]}(\mathbb{Z}_{\ell})\to\mathrm{GL}_{9[E:\mathbb{Q}]}(\mathbb{F}_{\ell})$ is $-\ell$, the surjective $\psi_{\ell}$ factors through $\bar{G}_{\ell}$. Then Lemma 3.8 implies that
(39) $\mathrm{JH}(\mathbb{Z}/(\ell-1)\mathbb{Z})\subset\mathrm{JH}(\prod_{i=1}^{k}\bar{\Gamma}_{\ell}^{\operatorname{ad}}).$
[…]
$[\Gamma_{\ell}^{\operatorname{ad}}:\alpha_{\ell}(\Gamma_{\ell}^{\operatorname{sc}})]\leq C$
for some constant $C$ depending only on $9[E:\mathbb{Q}]$. Thus, one obtains
(40) $[\bar{\Gamma}_{\ell}^{\operatorname{ad}}:\overline{\alpha_{\ell}(\Gamma_{\ell}^{\operatorname{sc}})}]\leq C$
where $\overline{\alpha_{\ell}(\Gamma_{\ell}^{\operatorname{sc}})}$ is the image of $\alpha_{\ell}(\Gamma_{\ell}^{\operatorname{sc}})$ in $\bar{\Gamma}_{\ell}^{\operatorname{ad}}$. Since $\bar{\Gamma}_{\ell}^{\operatorname{ad}}$ is hyperspecial maximal compact, it is equal to $\mathcal{H}_{\mathbb{Z}_{\ell}}(\mathbb{Z}_{\ell})$ for some semisimple group scheme $\mathcal{H}_{\mathbb{Z}_{\ell}}$ over $\mathbb{Z}_{\ell}$ whose generic fiber is $\mathbf{H}_{\ell}^{\operatorname{sc}}$. The special fiber $\mathcal{H}_{\mathbb{F}_{\ell}}$ is connected semisimple over $\mathbb{F}_{\ell}$ and of rank bounded by $9[E:\mathbb{Q}]$. Since the kernel of
$\Gamma_{\ell}^{\operatorname{sc}}=\mathcal{H}_{\mathbb{Z}_{\ell}}(\mathbb{Z}_{\ell})\twoheadrightarrow\mathcal{H}_{\mathbb{F}_{\ell}}(\mathbb{F}_{\ell})$
is $-\ell$ and $c_{\ell}(\mathcal{H}_{\mathbb{F}_{\ell}}(\mathbb{F}_{\ell}))$ is bounded by a constant $C^{\prime}$ depending only on $9[E:\mathbb{Q}]$. As $\Gamma_{\ell}^{\operatorname{sc}}$ surjects onto $\overline{\alpha_{\ell}(\Gamma_{\ell}^{\operatorname{sc}})}$, one obtains
(41) $c_{\ell}(\overline{\alpha_{\ell}(\Gamma_{\ell}^{\operatorname{sc}})})\leq C^{\prime}.$
Hence, (40), (41), and (ii) imply that
(42) $c_{\ell}(\prod_{i=1}^{k}\bar{\Gamma}_{\ell}^{\operatorname{ad}})\leq(CC^{\prime})^{k}\leq(CC^{\prime})^{[L:\mathbb{Q}]}.$
Therefore, (39) cannot hold (and thus $\zeta_{\ell}\notin\Omega_{\ell}$) if $\ell-1>(CC^{\prime})^{[L:\mathbb{Q}]}$. It follows that (34) holds for all $\ell$ belonging to some density one set $\mathcal{L}^{\prime\prime\prime}$.
We are done by taking $\mathcal{L}$ to be the intersection $\mathcal{L}^{\prime\prime}\cap\mathcal{L}^{\prime\prime\prime}$. ∎
Acknowledgments
G. Böckle acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124. C.-Y. Hui was partially supported by NSFC (no. 12222120) and a Humboldt Research Fellowship.